### v2

# Free Extreme points span free spectrahedrops and generalized free spectrahedra

Abstract.
Matrix convexity generalizes convexity to the dimension free setting and has connections to many mathematical and applied pursuits including operator theory, quantum information, noncommutative optimization, and linear control systems. In the setting of classical convex sets, extreme points are central objects which exhibit many important properties. For example, the well-known Minkowski theorem shows that any element of a closed bounded convex set can be expressed as a convex combination of extreme points. Extreme points are also of great interest in the dimension free setting of matrix convex sets; however, here the situation requires more nuance.
In the dimension free setting, there are many different types of extreme points. Of particular importance are free extreme points, a highly restricted type of extreme point that is closely connected to the dilation theoretic Arveson boundary. If free extreme points span a matrix convex set through matrix convex combinations, then they satisfy a strong notion of minimality in doing so. However, not all closed bounded matrix convex sets even have free extreme points. Thus, a major goal is to determine which matrix convex sets are spanned by their free extreme points.
Building on a recent work of J. W. Helton and the author which shows that free spectrahedra, i.e., dimension free solution sets to linear matrix inequalities, are spanned by their free extreme points, we establish two additional classes of matrix convex sets which are the matrix convex hull of their free extreme points. Namely, we show that closed bounded free spectrahedrops, i.e, closed bounded projections of free spectrahedra, are the span of their free extreme points. Furthermore, we show that if one considers linear operator inequalities that have compact operator defining tuples, then the resulting “generalized” free spectrahedra are spanned by their free extreme points.
Key words and phrases:
matrix convex set, extreme point, dilation theory, linear matrix inequality (LMI), spectrahedron, spectrahedrop, free Minkowski theorem, Arveson boundary, real algebraic geometry
2010 Mathematics Subject Classification:
Primary 47L07. Secondary 46L07, 90C22
1. Introduction
A linear matrix inequality (LMI) is an inequality of the form
$L_{A}(x):=I-A_{1}x_{1}-\dots A_{g}x_{g}\succeq 0$
where the $A_{1},\dots,A_{g}$ are $d\times d$ symmetric matrices and the $x_{1},\dots,x_{g}$ are real numbers. The solution set of such an LMI, i.e., the set of $x\in\mathbb{R}^{g}$ such that $L_{A}(x)\succeq 0$, is a convex set called a spectrahedron. Spectrahedra have connections to many areas including convex analysis, optimization, and linear systems control.
Linear matrix inequalities easily extend to the case where each $X_{\ell}$ is a $n\times n$ self-adjoint matrix whose product with $A_{i}$ is the Kronecker product. That is,
$L_{A}(X):=I_{dn}-A_{1}\otimes X_{1}-\dots-A_{g}\otimes X_{g}\succeq 0.$
A free spectrahedron is the set of all $g$-tuples of symmetric matrices (of any size) such that $L_{A}(X)\succeq 0$. The term “free” here refers to both the fact that the linear matrix inequality $L_{A}(X)\succeq 0$ is defined independent of the size of the matrices $X_{\ell}$, and the fact that the corresponding free spectrahedron contains matrix tuples of all sizes $n\times n$.
Free spectrahedra are prototypical examples of sets that exhibit a dimension-free type of convexity. Namely, free spectrahedra are matrix convex, i.e., are closed under matrix convex combinations where contraction matrices summing to the identity play the role of the convex coefficients. An important feature of matrix convex combinations is that they allow for combinations of matrix tuples of different sizes. This means that the geometry of any individual level of a matrix convex set is connected to that of all other levels of the set.
Another important example of matrix convex sets are projections of free spectrahedra [35]. Thanks to their association to linear matrix inequalities, these so called free spectrahedrops are more tractable than general matrix convex sets while also being more general than free spectrahedra. Free spectrahedrops and free spectrahedra have for example found use in optimization settings where one would like to determine if a given convex set is “LMI representable”, i.e., if it is the projection of a spectrahedron [34]. In addition, free spectrahedrops and free spectrahedra come up in problems related to spectrahedral inclusion [33, 12, 28, 56, 36] and in linear control engineering [37].
Mirroring the role of extreme points in classical convex sets, extreme points play a key role in the understanding of matrix convex sets [25, 29, 47, 23, 15]. However, in the dimension free setting of matrix convexity, there are many types of extreme point. A central goal in the study of matrix convex sets is to determine the most restricted type of extreme point that can recover any element of a given closed bounded matrix convex set through matrix convex combinations. That is, one searches for the strongest possible extension of the classical Minkowski theorem to the dimension free setting.
Of particular note in this pursuit are matrix extreme points [55, 26, 27] and free extreme points [41]. Matrix extreme points are known to span general closed bounded matrix convex sets through matrix convex combinations [42, 32]; however, they are not necessarily a minimal spanning set. Free extreme points, on the other hand, are more restricted than matrix extreme points [25, 21]. Stated informally, a free extreme point of a matrix convex set is an element that can only be expressed via trivial matrix convex combinations of other elements. Free extreme points are of great interest due both to their close connection [25] to the dilation theoretic Arveson boundary [3, 5] and to the fact that they necessarily form a minimal spanning set if they do span. The short coming of free extreme points is that, when restricting to finite dimensions, they can fail to span a given closed bounded matrix convex set. In fact, there are closed bounded matrix convex sets which have no (finite dimensional) free extreme points at all [22]. Thus an important question in the pursuit of a dimension free Minkowski theorem is “which matrix convex sets are the matrix convex hull of their free extreme points”.
1.1. Summary of Contributions
Recently, [24] showed that in the case of bounded free spectrahedra, free extreme points span. Furthermore, [24] showed that there is a tight dimension bound on the sum of sizes of free extreme points needed to express an element of a bounded free spectrahedron as a matrix convex combination of free extreme points. The main contribution of the present article is an extension of this result to two additional classes of matrix convex sets. Namely we show that closed bounded free spectrahedrops and closed bounded “generalized free spectrahedra”, i.e., closed bounded solution sets to linear operator inequalities with compact defining tuples, are the matrix convex hull of their free extreme points. An informal statement of this result is as follows.
Let $K$ be a closed bounded free spectrahedrop or a closed bounded generalized free spectrahedron over $\mathbb{F}=\mathbb{R}$ or $\mathbb{F}=\mathbb{C}$ and assume $K$ is closed under complex conjugation. If $X\in K$ is $g$-tuple of self-adjoint $n\times n$ matrices, then $X$ is a finite matrix convex combination of free extreme points of $K$ whose sum of sizes is bounded above by $n(g+1)$ if $\mathbb{F}=\mathbb{R}$ and by $2n(g+1)$ if $\mathbb{F}=\mathbb{C}$.
We point the reader to the upcoming Theorem 2.4 for a formal statement of our main result. Also see Sections 2.2 and 2.3 for formal definitions of generalized free spectrahedra and free spectrahedrops.
The proof of Theorem 2.4 follows the same approach used in [24] which itself was inspired by works such as [1, 20, 13]. Namely, we first work over $\mathbb{R}$ and show the existence of a special type of dilation called a maximal $1$-dilation. We then show that constructing finite sequence of at most $ng$ maximal $1$-dilations of $X$ results in a so called Arveson boundary point of $K$. Determining the irreducible components of the resulting Arveson boundary point yields an expression of $X$ as a matrix convex combination of free extreme points of $K$. An additional key ingredient which is need when $K$ is a free spectrahedrop is [19, Theorem 3.2] which allows us to express a free spectrahedrop as an intersection of free spectrahedra in a particularly well-behaved manner. A self contained version of this result which has been adapted to our setting is presented as Theorem 2.1.
The extension to the complex setting is achieved by a new result, Theorem 2.6, which shows that for matrix sets closed under complex conjugation, real free extreme points span the real part of the set if an only if the set is spanned by its free extreme points over the complexes. Similar to the methodology used in [24], this is accomplished using the standard embedding of tuples of $n\times n$ complex self-adjoint matrices to tuples of $2n\times 2n$ real symmetric matrices together with showing that, for sets closed under complex conjugation, a free extreme point over the reals is also a free extreme point over the complexes. A key feature of Theorem 2.6 is that it is proved for general matrix convex sets closed under complex conjugation, while [24] only treated the case of free spectrahedra.
Finally, we give a new necessary condition for a closed bounded matrix convex set to be a free spectrahedrop, see Corollary 2.7. Namely, if $K$ is closed bounded free spectrahedrop and $K$ is closed under complex conjugation, then both $K$ and its free polar dual are the matrix convex hull of their free extreme points. This corollary provides new insight into the open question of which matrix convex sets are free spectrahedrops. In Proposition 6.2, we show the condition is nontrivial and use it to show the existence of generalized free spectrahdra that are not free spectrahedrops.
1.2. Related works
The study of extreme points of matrix convex sets goes back Arveson who, (in the language of completely positive maps on operator systems) conjectured that if one extends to infinite dimensional levels, then infinite dimensional free extreme points span [3, 4]. This infinite dimensional question was studied by a number of authors [31, 1, 45, 20, 5, 41, 46] until it was finally settled in 2015 by Davidson and Kennedy [13]. Work on extreme points in this setting has since continued, e.g., see [14] which develops Choquet theory for the infinite dimensional analog of matrix convex sets.
Matrix convexity is also closely related to the rapidly growing area of free analysis [7, 51, 17, 53, 39, 2, 6, 38]. Here the goal is to extend classical geometric and function theoretic results to the noncommutative setting where one considers functions whose inputs are $g$-tuples of matrices or operators. This study was largely pushed by Voiculescu’s introduction of free probability which has since been used to great effect in random matrix theory [52, 44]. Other closely related topics include noncommutative optimization [18, 50, 10, 43, 54] and quantum information and games [9, 8, 30, 16, 49].
1.3. Reader’s guide
Section 2 introduces our definitions and notation and gives a formal statement of our main results, Theorem 2.4, Theorem 2.6, and Corollary 2.7. Section 3 introduces the notion of a maximal $1$-dilation for free spectrahedrops and shows that maximal $1$-dilations in a bounded real free spectrahedrop reduce the dimension of the dilation subspace. Section 4 is the equivalent of Section 3 in the case of real generalized free spectrahedra. Section 5 handles the extension from reals to complexes. Finally, Section 6 discusses the relationship between generalized free spectrahedra and free spectrahedrops and gives an example of a generalized free spectrahedron that is not a free spectrahedrop.
2. Definitions, notation, and main results
Throughout the article we let ${\mathcal{H}}$ denote a Hilbert space over $\mathbb{F}=\mathbb{R}$ or $\mathbb{F}=\mathbb{C}$. We let $B({\mathcal{H}})$ and $SA({\mathcal{H}})$ respectively denote the sets of bounded operators on ${\mathcal{H}}$ and bounded self-adjoint operators on ${\mathcal{H}}$. Additionally let $SA({\mathcal{H}})^{g}$ denote the set of $g$-tuples of the form $X=(X_{1},\dots,X_{g})$ where each $X_{\ell}$ is a bounded self-adjoint operator on ${\mathcal{H}}$. Say tuples $X,Y\in SA({\mathcal{H}})^{g}$ are unitarily equivalent if there exists a unitary $U\in B({\mathcal{H}})$ such that
$U^{*}XU=(U^{*}X_{1}U,\dots,U^{*}X_{g}U)=(Y_{1},\dots,Y_{g})=Y.$
A bounded operator $B\in B({\mathcal{H}})$ is positive semidefinite if it is self-adjoint and for any $v\in{\mathcal{H}}$, one has $\langle Bv,v\rangle\geq 0$. In the case that $B$ is a compact self-adjoint operator, this is equivalent to all of $B$’s eigenvalues being nonnegative. Given two bounded self-adjoint operators $B_{1},B_{2}\in SA({\mathcal{H}})$ let $B_{1}\succeq B_{2}$ denote that $B_{1}-B_{2}$ is positive semidefinite.
While general Hilbert spaces do play a role in this article, our main interest is matrix convex sets which are restricted to finite dimensions. Thus it is convenient so also have notation for matrix spaces. We let $M_{m\times n}(\mathbb{F})^{g}$ denote the set of $g$-tuples of $m\times n$ matrices with entries in $\mathbb{F}$, and let $M_{n}(\mathbb{F})^{g}=M_{n\times n}(\mathbb{F})^{g}$. Additionally, $SM_{n}(\mathbb{F})^{g}$ is the set of all $g$-tuples of real self-adjoint (symmetric) $n\times n$ matrices. Also, we set $SM(\mathbb{F})^{g}=\cup_{n}SM_{n}(\mathbb{F})^{g}$. Given a subset $K\subseteq SM(\mathbb{F})^{g}$, we let $K(n)$ denote the set
$K(n):=K\cap SM_{n}(\mathbb{F})^{g}.$
That is $K(n)$ is set of $g$-tuples of $n\times n$ matrices that are elements of $K$. The set $K(n)$ is called the $n$th level of $K$.
For clarity we emphasize that, even when we restrict to real settings, we typically use the term self-adjoint rather than symmetric to provide consistency in the terminology throughout the article. Similarly, to be consistent with this terminology choice, we typically use $B^{*}$ rather than $B^{T}$ to denote the transpose of a matrix $B$.
2.1. Linear operator inequalities
Given a $g$-tuple
$A=(A_{1},\dots,A_{g})\in SA({\mathcal{H}})^{g}$
of self-adjoint operators on ${\mathcal{H}}$, a monic linear pencil is a sum of the form
$L_{A}(x)=I_{\mathcal{H}}-A_{1}x_{1}-A_{2}x_{2}-\dots-A_{g}x_{g}.$
Given a tuple $X\in SM_{n}(\mathbb{F})^{g}$, the evaluation of $L_{A}$ at $X$ is
$L_{A}(X)=I_{{\mathcal{H}}}\otimes I_{n}-A_{1}\otimes X_{1}-A_{2}\otimes X_{2}-\dots-A_{g}\otimes X_{g}$
where $\otimes$ denotes the Kronecker tensor product. Additionally, let $\Lambda_{A}(X)$ denote the homogeneous linear part of $L_{A}(X)$. That is,
$\Lambda_{A}(X)=A_{1}\otimes X_{1}+A_{2}\otimes X_{2}+\dots+A_{g}\otimes X_{g},$
The inequality
$L_{A}(X)\succeq 0,$
is called a linear operator inequality. Moreover, if ${\mathcal{H}}$ is finite dimension, then the inequality $L_{A}(X)\succeq 0$ is called a linear matrix inequality.
2.2. Free spectrahedra
Given a $g$-tuple $A\in SA({\mathcal{H}})^{g}$ and a positive integer $n$ define the set ${\mathcal{D}}_{A}(n)\subseteq SM_{n}(\mathbb{F})^{g}$ by
$\mathcal{D}_{A}(n):=\{X\in SM_{n}(\mathbb{F})^{g}:L_{A}(X)\succeq 0\}.$
That is, ${\mathcal{D}}_{A}(n)$ is the set of all $g$-tuples of $n\times n$ self-adjoint matrices $X$ such that the evaluation $L_{A}(X)$ is positive semidefinite. Additionally define the set ${\mathcal{D}}_{A}\subseteq SM(\mathbb{F})^{g}$ to be the union over all $n$ of the the sets ${\mathcal{D}}_{A}(n)$. That is,
$\mathcal{D}_{A}:=\bigcup_{n=1}^{\infty}\mathcal{D}_{A}(n).$
If ${\mathcal{H}}$ is finite dimensional, then ${\mathcal{D}}_{A}$ is called a free spectrahedron and ${\mathcal{D}}_{A}(n)$ is the free spectrahedron ${\mathcal{D}}_{A}$ at level $n$ [25] . In this article, we extend to considering the case where ${\mathcal{H}}$ is a separable infinite dimensional Hilbert space and $A\in SA({\mathcal{H}})^{g}$ is a compact operator tuple. In this case we call ${\mathcal{D}}_{A}$ a generalized free spectrahedron and we call ${\mathcal{D}}_{A}(n)$ a generalized free spectrahedron at level $n$. That is, a generalized free spectrahedron ${\mathcal{D}}_{A}$ is the solution set of the linear operator inequality
$L_{A}(X)\succeq 0$
where $A$ is a tuple of self-adjoint compact operators on ${\mathcal{H}}$.
2.3. Free spectrahedrops
Central to this article are (coordinate) projections of free spectrahedra, i.e., free spectrahedrops. Given a finite dimensional Hilbert space ${\mathcal{H}}$ and a $g$-tuple $A\in SA({\mathcal{H}})^{g}$ and an $h$-tuple $B\in SA({\mathcal{H}})^{h}$ of self-adjoint operators on ${\mathcal{H}}$, let $L_{(A,B)}(x,y)$ denote the monic linear pencil
$L_{(A,B)}(x,y)=I_{{\mathcal{H}}}-A_{1}x_{1}-\dots-A_{g}x_{g}-B_{1}y_{1}-\dots-B_{h}y_{y}.$
Also let ${\mathcal{D}}_{(A,B)}$ denote the free spectrahedron
${\mathcal{D}}_{(A,B)}=\{(X,Y)\in SM(\mathbb{F})^{g+h}:L_{(A,B)}(X,Y)\succeq 0\}.$
Define the set ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ by
${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}:=\{X\in{SM(\mathbb{F})^{g}}:\ \mathrm{There\ exists\ }Y\in SM(\mathbb{R})^{h}\ \mathrm{such\ that\ }(X,Y)\in{\mathcal{D}}_{(A,B)}\}.$
The set ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a called a free spectrahedrop.
Let $K$ be a free spectrahedrop or generalized free spectrahedron. Say $K$ is bounded if there is some real number $C$ so that
$CI_{n}-\sum_{i=1}^{g}X_{i}^{2}\succeq 0$
for all $X=(X_{1},X_{2},\dots,X_{g})\in K(n)$ and all positive integers $n$. It is not difficult to show that such a set $K$ is bounded if and only if $K(1)$ is bounded.
In addition, we say $K$ is closed if $K(n)$ is closed for all $n$. We note that the generalized free spectrahedra considered in this article are, by definition, closed in this sense. However, free spectrahedrops are not necessarily closed, as the projection of an (unbounded) closed set can fail to be closed.
2.4. Drescher, Netzer, and Thom vs Free spectrahedrops
Critical to the proof of our main result for free spectrahedrops is a result of Drescher, Netzer, and Thom [19], which affords us a clean membership test for elements of a free spectrahedrop. To present this result, we introduce the following notation.
Given $d$-dimensional Hilbert space ${\mathcal{H}}$ and an $h$-tuple $B\in SA({\mathcal{H}})^{h}$ of self-adjoint operators on ${\mathcal{H}}$, let $B^{\oplus n}$ denote the direct sum of $B$ with itself $n$ times. Additionally, define the sets ${\mathcal{Z}}_{B}(n)$ and ${\mathcal{I}}_{B}(n,r)$ and ${\mathcal{I}}_{B}(n)$ by
$\begin{array}[]{rclcl}{\mathcal{Z}}_{B}(n)&=&\{W\in B({\mathcal{H}}^{nd},\mathbb{F}^{n}):\ W^{*}(B^{\oplus nd})W=0_{n\times n}\}\\ {\mathcal{I}}_{B}(n,r)&=&\{W\in B({\mathcal{H}}^{nd},\mathbb{F}^{r}):\ W^{*}(B^{\oplus nd})W=0_{r\times r}\mathrm{\ and\ }W^{*}W=I_{r}\}\\ {\mathcal{I}}_{B}(n)&=&\cup_{r\leq n}\ {\mathcal{I}}_{B}(n,r).\end{array}$
We will frequently make use of the fact that, for fixed $n$ and $r$, the set ${\mathcal{I}}_{B}(n,r)$ is compact.
Theorem 2.1 ([19, Theorem 3.2, Proposition 3.3]).
Let $A\in{SM_{d}(\mathbb{F})^{g}}$ and $B\in{SM_{d}(\mathbb{F})^{h}}$ such that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded free spectrahedrop and $X\in{SM_{n}(\mathbb{F})^{g}}$. Then we have the following
  1. (1)
    $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$ if and only if for all $W\in{\mathcal{Z}}_{B}(n)$ one has
    (2.1) $W^{*}W\otimes I_{n}-{\Lambda_{W^{*}\left(A^{\oplus nd}\right)W}}(X)\succeq 0.$
  2. (2)
    $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$ if an only if for all $W\in{\mathcal{I}}_{B}(n)$ one has
    (2.2) ${L_{W^{*}\left(A^{\oplus nd}\right)W}}(X)\succeq 0.$
  3. (3)
    If $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$, then for any $m\in{\mathbb{N}}$ and any $W\in{\mathcal{I}}_{B}(m)$, one has
    ${L_{W^{*}\left(A^{\oplus md}\right)W}}(X)\succeq 0.$
Proof.
The proof of Item (1) is essentially the same as that of [19, Theorem 3.2], with minor modifications to handle the case $\mathbb{F}=\mathbb{R}$. The proof of Item (2) follows from the same approach used to prove [19, Proposition 3.3]. Finally, the proof of Item (3) quickly follows from Items (1) and (2) together a straightforward argument using techniques that are routine in the study of matrix convex sets. Since this theorem plays a key role in our upcoming proofs, and since the setting and statement here is slightly different than in [19], a self-contained version of the proof is given in Appendix 7. ∎
2.5. Matrix Convex Sets
Given a finite collection of $g$-tuples $\left\{X^{i}\right\}_{i=1}^{\ell}$ with $X^{i}\in SM_{n_{i}}(\mathbb{F})^{g}$ for each $i=1,2,\dots,\ell$, a matrix convex combination of $\left\{X^{i}\right\}_{i=1}^{\ell}$ is a sum of the form
$\sum_{i=1}^{\ell}V_{i}^{*}X^{i}V_{i}\qquad\text{with}\qquad\sum_{i=1}^{\ell}V_{i}^{*}V_{i}=I_{n}$
where $V_{i}\in M_{n_{i}\times n}(\mathbb{F})$ and
$V_{i}^{*}X^{i}V_{i}=\left(V_{i}^{*}X_{1}^{i}V_{i},V_{i}^{*}X_{2}^{i}V_{i},\dots,V_{i}^{*}X_{g}^{i}V_{i}\right)\in SM_{n}(\mathbb{R})^{g}$
for all $i=1,2,\dots,\ell$. A key feature of matrix convex combinations is that the tuples $X^{i}$ need not be the same size.
A set $K\subseteq SM(\mathbb{F})^{g}$ is matrix convex if it is closed under matrix convex combinations and the matrix convex hull of $K$, denoted $\mathrm{co}^{\mathrm{mat}}(K)$, is the set of all matrix convex combinations of the elements of $K$. Equivalently, $K$ is matrix convex if and only if $K=\mathrm{co}^{\mathrm{mat}}(K)$. It is straightforward to show that generalized free spectrahedra and free spectrahedrops are matrix convex. Say a matrix convex set $K$ is closed under complex conjugation if $X=(X_{1},\dots,X_{g})\in K$ implies $\overline{X}=(\overline{X}_{1},\dots,\overline{X}_{g})\in K$.
2.6. Free extreme points of matrix convex sets
As previously mentioned, matrix convex sets have many different types of extreme points, e.g., see [25, 42, 40]. In this article, we restrict our attention to free extreme points. Given a set $K\subseteq SM(\mathbb{F})^{g}$, we say a point $X\in K(n)$ is a free extreme point of $K$ if whenever $X$ is written as a matrix convex combination
$X=\sum_{i=1}^{\ell}V_{i}^{*}X^{i}V_{i}\qquad\text{with}\quad\sum_{i=1}^{\ell}V_{i}^{*}V_{i}=I_{n},$
of points $X^{i}\in K$ with $V_{i}\neq 0$ for each $i$, then for all $i=1,2,\dots,\ell,$ either $V_{i}\in M_{n}(\mathbb{F})$ and $X$ is unitarily equivalent to $X^{i}$ or $V_{i}\in M_{n_{i}\times n}(\mathbb{F})$ where $n_{i}>n$ and there exists $Z^{i}\in K$ such that $X\oplus Z^{i}$ is unitarily equivalent to $X^{i}$. Intuitively, a tuple $X$ is a free extreme point of $K$ if it cannot be written as a nontrivial matrix convex combination of any collection of elements of $K$. We let $\partial^{\mathrm{free}}{K}$ denote the set of all free extreme points of $K$.
2.7. Free extreme points, dilations, and the Arveson boundary
We next discuss the connection between free extreme points and the dilation theoretic Arveson boundary. To do this, we must first introduce the notion of an irreducible tuple of matrices.
Given a matrix $M\in M_{n}(\mathbb{F})$, a subspace $N\subseteq\mathbb{F}^{n}$ is a reducing subspace if both $N$ and $N^{\perp}$ are invariant subspaces of $M$, A tuple $X\in SM_{n}(\mathbb{R})^{g}$ is irreducible (over $\mathbb{F}$) if the matrices $X_{1},\dots,X_{g}$ have no common reducing subspaces in $\mathbb{F}^{n}$; a tuple is reducible (over $\mathbb{F}$) if it is not irreducible. For the remainder of the article, we drop the distinction “over $\mathbb{F}$” when referring to irreducible tuples. However, we warn the reader that a tuple may be reducible over $\mathbb{C}$ even if it is irreducible over $\mathbb{R}$. Thus irreducibility over $\mathbb{R}$ is not equivalent to other well-known definitions of irreducibility which are instead equivalent to irreducibility over $\mathbb{C}$.
2.7.1. Dilations
Let $K\subseteq{SM(\mathbb{F})^{g}}$ be a matrix convex set and let $X\in K(n)$. If there exists a positive integer $\ell\in{\mathbb{N}}$ and $g$-tuples $\beta\in M_{n\times\ell}(\mathbb{F})^{g}$ and $\gamma\in{SM_{\ell}(\mathbb{F})^{g}}$ such that
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}=\left(\begin{pmatrix}X_{1}&\beta_{1}\\ \beta^{*}_{1}&\gamma_{1}\end{pmatrix},\cdots,\begin{pmatrix}X_{g}&\beta_{g}\\ \beta^{*}_{g}&\gamma_{g}\end{pmatrix}\right)\in K,$
then we say $Y$ is an $\ell$-dilation of $X$. The tuple $Y$ is a trivial dilation of $X$ if $\beta=0$. A key connection between matrix convex combinations and dilations is the following. If $Y$ is a dilation of $X$ and $V^{*}=\begin{pmatrix}I_{n}&0\end{pmatrix},$ then $X=V^{*}YV$ with $V^{*}V=I_{n}$. That is, if $Y$ is a dilation of $X$, then $X$ can be expressed as a matrix convex combination of $Y$. We note that this matrix convex combination is non-trivial if the dilation itself is non-trivial.
Given a matrix convex set $K$ and an element $X\in K$, we define the dilation subspace of $K$ at $X$, denoted ${\mathfrak{D}_{X}(K)}$, to be
${\mathfrak{D}_{X}(K)}=\mathrm{span}\left(\left\{\beta\in M_{n\times 1}(\mathbb{F})^{g}:\mathrm{there\ exists\ a\ }\gamma\in\mathbb{R}^{g}\mathrm{\ s.t.\ }\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K\right\}\right).$
The dilation subspace obtains its name as, up to scaling, it gives the space of viable dilations of an element of a matrix convex set.
Lemma 2.2.
Let $K\subseteq SM(\mathbb{F})^{g}$ be a matrix convex set and let $X\in K$. If $\beta\in{\mathfrak{D}_{X}(K)}$, then there exists a real constant $c>0$ and a tuple $\gamma\in\mathbb{R}^{g}$ such that
$\begin{pmatrix}X&c\beta\\ c\beta^{*}&\gamma\end{pmatrix}\in K.$
Proof.
If $\beta=0\in M_{n\times 1}(\mathbb{F})^{g}$, the proof is trivial, so assume $\beta\neq 0$. From the definition of the dilation subspace, the tuple $\beta$ can be expressed as a linear combination $\beta=\sum_{i=1}^{k}\alpha_{i}\beta^{i}$ of tuples $\beta^{i}\in M_{n\times 1}(\mathbb{F})^{g}$ where for each $i$, there exists a $\gamma^{i}\in\mathbb{R}^{g}$ such that
$\begin{pmatrix}X&\beta^{i}\\ (\beta^{i})^{*}&\gamma^{i}\end{pmatrix}\in K.$
We first argue that the $\alpha_{i}$ can be taken to be positive real numbers. To this end, for each $i$ write $\alpha_{i}=|\alpha_{i}|e^{i\theta}$ in polar coordinates and observe that
$\begin{pmatrix}1&0\\ 0&e^{-i\theta}\end{pmatrix}\begin{pmatrix}X&\beta^{i}\\ (\beta^{i})^{*}&\gamma^{i}\end{pmatrix}\begin{pmatrix}1&0\\ 0&e^{i\theta}\end{pmatrix}=\begin{pmatrix}X&e^{i\theta}\beta^{i}\\ (e^{i\theta}\beta^{i})^{*}&\gamma^{i}\end{pmatrix}\in K.$
Setting $\tilde{\beta}^{i}=e^{i\theta}\beta^{i}$, we obtain $\beta=\sum_{i=1}^{k}|\alpha_{i}|\tilde{\beta}^{i}$. Therefore, we can without loss of generality assume $\alpha_{i}>0$ as claimed.
Now, define $c\in\mathbb{R}$ by $1/c=\sum_{i=1}^{k}\alpha_{i}$. Then $\sum_{i=1}^{k}c\alpha_{i}=1$, so the convex combination
$\sum_{i=1}^{k}(c\alpha_{i})\begin{pmatrix}X&\beta^{i}\\ (\beta^{i})^{*}&\gamma^{i}\end{pmatrix}=\begin{pmatrix}X&\sum_{i=1}^{k}c\alpha_{i}\beta^{i}\\ \sum_{i=1}^{k}c\alpha_{i}(\beta^{i})^{*}&\sum_{i=1}^{k}c\alpha_{i}\gamma^{i}\end{pmatrix}=\begin{pmatrix}X&c\beta\\ c\beta^{*}&\sum_{i=1}^{k}c\alpha_{i}\gamma^{i}\end{pmatrix}.$
is an element of $K$, which completes the proof. ∎
In the case that $K$ is a free spectrahedron, the definition of the dilation subspace given here is equivalent to the definition given in [24], though the presentation is different, see [24, Lemma 2.1]. We present the definition in this form so that the definition applies to general matrix convex sets. This also allows us to use a single definition for both free spectrahedrops and generalized free spectrahedra. See Lemma 3.1 and Lemma 4.4 for further discussion.
2.7.2. Arveson extreme points
A tuple $X\in K$ is an Arveson extreme point of $K$ if $K$ does not contain a nontrivial dilation of $X$. More precisely, $X\in K$ is an Arveson extreme point of $K$ if and only if, if
(2.3) $\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K(n+\ell),$
for some tuples $\beta\in M_{n\times\ell}(\mathbb{R})^{g}$ and $\gamma\in SM_{\ell}(\mathbb{R})^{g}$, then $\beta=0$. Equivalently, $X$ is an Arveson extreme point of $K$ if and only if $\dim{\mathfrak{D}_{X}(K)}=0$. If $Y$ is an Arveson extreme point of $K$ and $Y$ is an ($\ell$-)dilation of $X\in K$, then we will say $Y$ is an Arveson ($\ell$-)dilation of $X$.
The following theorem gives the connection between free and Arveson extreme points.
Theorem 2.3 ([25, Theorem 1.1 (3)]).
Let $K\subseteq SM(\mathbb{F})^{g}$ be a matrix convex set. Then $X$ is a free extreme point of $K$ if and only if $X$ is an Arveson extreme point of $K$ and is irreducible.
This theorem has the consequence that the irreducible components of an Arveson extreme point are all free extreme points. This means that expressing an element $X$ of a matrix convex set as a matrix convex combination of free extreme points can be accomplished by finding an Arveson dilation of $X$.
2.8. Main results
We are now in position to give a formal statement of our main spanning result for free spectrahedrops and generalized free spectrahedra.
Theorem 2.4.
Let $K\subseteq SM(\mathbb{F})^{g}$ be a closed bounded free spectrahedrop or a closed bounded generalized free spectrahedron and assume $K$ is closed under complex conjugation. Given an $X\in K(n)$, there exists an Arveson $\ell$-dilation $Y\in K(n+\ell)$ of $X$ where $\ell$ satisfies
(2.4) $\ell\leq n+2\dim{\mathfrak{D}_{X}(K)}\leq n+2ng.$
Furthermore, if $X\in SM_{n}(\mathbb{R})^{g}$ is real valued, then $\ell$ satisfies
(2.5) $\ell\leq\dim{\mathfrak{D}_{X}(K)}\leq ng.$
As a consequence, $X$ can be written as a matrix convex combination
$X=\sum_{i=1}^{k}V_{i}^{*}X^{i}V_{i}\qquad\mathrm{s.t.}\qquad\sum_{i=1}^{k}V_{i}^{*}V_{i}=I$
of free extreme points $X^{i}\in K(n_{i})$ of $K$ where $\sum_{i=1}^{k}n_{i}\leq n+\ell$. Thus, $K$ is the matrix convex hull of its free extreme points.
Remark 2.5.
Working over $\mathbb{C}$, we obtain $n+\ell\leq 2n(g+1)$ while over $\mathbb{R}$ we obtain $n+\ell\leq n(g+1)$. The extra factor of $2$ in the complex bound comes from embedding a tuple of $n\times n$ complex matrices into a tuple of $2n\times 2n$ real matrices.
Proof.
The proof over $\mathbb{R}$ of the first part of Theorem 2.4 quickly follows from Theorems 3.7 and 4.6. In particular, let $X\in K$, and set $m=\dim{\mathfrak{D}_{X}(K)}$. Assume that $X$ is not an Arveson extreme point of $K$, i.e., that $m>0$. Also note that $m\leq ng$ since ${\mathfrak{D}_{X}(K)}$ is a subspace of $M_{n\times 1}(\mathbb{R})^{g}$.
Using Theorem 3.7 when $K$ is a bounded free spectrahedrop and Theorem 4.6 when $K$ is a bounded generalized free spectrahedron shows that there exists an integer $\ell\leq m$ and a collection of tuples $\{Y^{i}\}_{i=0}^{\ell}\subseteq K$ such that the following hold:
  1. (1)
    $Y^{0}=X$.
  2. (2)
    For each $i=1,\dots,\ell-1$, the tuple $Y^{i+1}$ is a $1$-dilation of $Y^{i}$ and
    $\dim{\mathfrak{D}_{Y^{i+1}}(K)}<\dim{\mathfrak{D}_{Y^{i}}(K)}.$
  3. (3)
    $Y^{\ell}$ satisfies $\dim{\mathfrak{D}_{Y^{\ell}}(K)}=0$.
It follows from Lemmas 3.1 (3) and 4.4 (3) that $Y$ is an Arveson extreme point of $K$. Thus $Y^{\ell}\in K(n+\ell)$ is an Arveson $\ell$-dilation of $X$.
The proof over $\mathbb{R}$ of the second part of the theorem quickly follows the first part together with routine dilation theoretic arguments. In particular, decomposing the tuple $Y^{\ell}$ into its irreducible components yields the desired collection of free extreme points. See the proof of [24, Theorem 1.3] for a similar argument.
The proof of the theorem over $\mathbb{C}$ follows by combining the theorem over the reals together with the upcoming Theorem 2.6, Lemma 5.3, and Lemma 5.4 (4). More precisely, Theorem 2.6 shows that for a matrix convex set closed under complex conjugation, free extreme points spanning over the complexes reduces to real free extreme points spanning the real part of the set. Next, Lemma 5.3 shows that the real part of a free spectrahedrop/generalized free spectrahedron is indeed a real free spectrahedrop/generalized free spectrahedron, hence we can indeed apply the real spanning theorem. Finally, Lemma 5.4 (4) yields the upper bound presented in equation (2.4) on the sum of size of the dilation required. ∎
Our second main result shows that, for matrix convex sets that our closed under complex conjugation, free extreme points spanning over the complexes reduces to free extreme points spanning over the reals.
Theorem 2.6.
Let $K\subseteq SM(\mathbb{C})^{g}$ be a matrix convex set such that $K=\overline{K}$, and let $K^{\mathbb{R}}$ denote $K\cap SM(\mathbb{R})^{g}$. Then $K$ is the matrix convex hull of its free extreme points if and only if $K^{\mathbb{R}}$ is the matrix convex hull of its free extreme points. In notation,
$K=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K))\qquad\mathrm{if\ and\ only\ if}\qquad K^{\mathbb{R}}=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K^{\mathbb{R}})).$
In addition, given a tuple $X\in K$, write $X=X_{\mathbb{R}}+iX_{\mathbb{C}}\in K$ where $X_{\mathbb{R}}$ and $X_{\mathbb{C}}$ are the real and imaginary parts of $X$. If the tuple
$\begin{pmatrix}X_{\mathbb{R}}&X_{\mathbb{C}}\\ X_{\mathbb{C}}^{T}&X_{\mathbb{R}}\end{pmatrix}\in K^{\mathbb{R}}$
can be expressed as a matrix convex combination of free extreme points of $K^{\mathbb{R}}$ whose sum of sizes is at most $m$, then $X$ can also be expressed as a matrix convex combination of free extreme points of $K$ whose sum of sizes is at most $m$.
Proof.
See Section 5.1. ∎
Theorem 2.4 also has implications for another open question in the study of free spectrahedrops. Namely, is there a necessary and sufficient condition for a matrix convex set $K$ to be a closed bounded free spectrahedrop? From [35, Theorem 4.11], if $K\subseteq SM(\mathbb{F})^{g}$ is a closed bounded free spectraherop, then the polar dual of $K$ is also a closed bounded free spectrahedrop. Here the polar dual of $K$, denoted $K^{\circ}$ is defined by
$K^{\circ}=\{A\in SM(\mathbb{F})^{g}:L_{X}(A)\succeq 0\mathrm{\ for\ all\ }X\in K\}.$
Corollary 2.7.
If $K\subseteq SM(\mathbb{F})^{g}$ is a closed bounded free spectrahedrop that is closed under complex conjugation, then both $K$ and $K^{\circ}$ are the matrix convex hull of their free extreme points.
Proof.
Immediate from Theorem 2.4 together with [35, Theorem 4.11]. ∎
In the upcoming Theorem 6.1 and Proposition 6.2 we give an example of a closed bounded matrix convex set $K$ such that $K$ is the matrix convex hull of its free extreme points while $K^{\circ}$ is not. This illustrates that Corollary 2.7 provides nontrivial restrictions that a matrix convex set $K$ must satisfy to be a closed bounded free spectrahedrop.
2.8.1. Matrix convex sets that are not the matrix convex hull of their free extreme points
We conclude the section by discussing the limits of possible spanning theorems for free extreme points. We first note that if one extends to “noncommutative convex sets” which include infinite dimensional levels, then every compact matrix convex set is the closed “noncommutative convex hull” of its “nc extreme points”, which are an analogue of free extreme points in the infinite dimensional setting, see [14, Theorem 6.4.2]. However, it is natural to ask if what starts in finite dimensions stays in finite dimensions. To this end, there are several known examples of matrix convex sets that are not the matrix convex hull of their (finite dimensional) free extreme points.
For example, [42, Example 6.30] illustrates how one can arrive at a matrix convex set that has no free extreme points by considering generators of the Cuntz algebra. In fact, the resulting matrix convex set is a free spectrahedron over $\mathbb{C}$ that is not closed under complex conjugation [46, Remark 2.5]. This illustrates that the assumption that $K$ is closed under complex conjugation in Theorem 2.4 is necessary. Examples of real compact matrix convex sets that have no free extreme points are also known and are discussed in [22]. These example are essentially obtained by taking the matrix convex hull of an irreducible tuple of compact operators on a real Hilbert space.
To the author’s knowledge, all known examples of compact matrix convex sets that are closed under complex conjugation and are the matrix convex hull (without taking closure) of their (finite dimensional) free extreme points are either free spectrahedrops or generalized free spectrahedra. While the author does not expect that this is a necessary condition for spanning, it may give an interesting direction for future work.
3. Arveson extreme points and real free spectrahedrops
This section introduces maximal $1$-dilations for real free spectrahedrops and presents Theorem 3.7 which shows that maximal $1$-dilations in a real free spectrahedrop reduce the dimension of the dilation subspace. For emphasis, throughout the section we take $\mathbb{F}=\mathbb{R}$. Additionally, since we only consider finite dimensional spaces ${\mathcal{H}}$ when working with free spectrahedrops, throughout the section we set ${\mathcal{H}}=\mathbb{R}^{d}$.
We begin by establishing an alternative definition of the dilation subspace for real free spectrahedrops. Let $A\in SM_{d}(\mathbb{R})^{g}$ and $B\in SM_{d}(\mathbb{R})^{h}$ such that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded real free spectrahedrop. Given a tuple $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$, define the subspace ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\subseteq M_{n\times 1}(\mathbb{R})^{g}$ by
(3.1) ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}=\cap_{W\in{\mathcal{I}}_{B}(n+1)}\{\beta\in M_{n\times 1}(\mathbb{R})^{g}:\ \ker{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)\subseteq\ker{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})\}.$
Note that ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ is indeed a subspace of $M_{n\times 1}(\mathbb{R})^{g}$ since the set
(3.2) $\{\beta\in M_{n\times 1}(\mathbb{R})^{g}:\ \ker{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)\subseteq\ker{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})\}$
is a subspace of $M_{n\times 1}(\mathbb{R})^{g}$ for each fixed $W\in{\mathcal{I}}_{B}(n+1)$.
The following lemma shows that ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ is equal to the dilation subspace of ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ at $X$ and explains the connection between ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ and dilations of $X$.
Lemma 3.1.
Let ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}\subseteq SM(\mathbb{R})^{g}$ be a closed bounded real free spectrahedrop and let $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$.
  1. (1)
    If $\beta\in M_{n\times 1}(\mathbb{R})^{g}$ and
    $Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$
    is a $1$-dilation of $X$, then $\beta\in{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$.
  2. (2)
    Let $\beta\in M_{n\times 1}(\mathbb{R})^{g}$. Then $\beta\in{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ if and only if there is a tuple $\gamma\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(1)$ and a real number $c_{\gamma}>0$ such that
    $\begin{pmatrix}X&c_{\gamma}\beta\\ c_{\gamma}\beta^{*}&\gamma\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1).$
    In particular, one may take $\gamma=0\in\mathbb{R}^{g}$.
  3. (3)
    One has
    ${\mathfrak{D}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}={{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
    As a consequence, $X$ is an Arveson extreme point of ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ if and only if
    $\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}=0.$
Proof.
Item (1) follows from considering the Schur complement of ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)$ for a dilation
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$
of $X$. Indeed, multiplying ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)$ by permutation matrices, sometimes called canonical shuffles, see [48, Chapter 8], shows
(3.3) ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)\succeq 0\quad\mathrm{if\ and\ only\ if}\quad\begin{pmatrix}{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)&{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta)\\ {\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})&{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(\gamma)\end{pmatrix}\succeq 0.$
Taking the appropriate Schur complement then shows that ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)\succeq 0$ if and only if
(3.4) ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(\gamma)\succeq 0,$
$\mathrm{\ and\ }$ ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)-{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta){L_{W^{*}\left(A^{\oplus n+1}\right)W}}(\gamma)^{\dagger}{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})\succeq 0,$
where $\dagger$ denotes the Moore-Penrose pseudoinverse. It follows that if ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)\succeq 0$, then
$\ker{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)\subseteq\ker{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*}).$
Using Theorem 2.1 (2) shows that $Y\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$ if and only if ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(Y)\succeq 0$ for all $W\in{\mathcal{I}}_{B}(n+1)$. Combining this with the above, we conclude that if $Y\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$, then
$\ker{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)\subseteq\ker{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})$
for all $W\in{\mathcal{I}}_{B}(n+1)$.
We now prove Item (2). We first show that for each fixed $W\in{\mathcal{I}}_{B}(n+1)$ there exists some constant $c_{W}>0$ such that
(3.5) ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&c_{W}\beta\\ c_{W}\beta^{*}&0\end{pmatrix}\succeq 0.$
Note that ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(0)=I$, so similar to before using the Schur complement shows the above inequality holds if and only if
(3.6) ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)-c_{W}^{2}{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta){\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*})\succeq 0.$
If $\beta\in{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ then
$\ker{L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)\subseteq\ker{\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta){\Lambda_{W^{*}\left(A^{\oplus n+1}\right)W}}(\beta^{*}),$
so picking $c_{W}>0$ small enough so that $\|c_{W}^{2}\Lambda_{A}(\beta)\Lambda_{A}(\beta^{*})\|_{2}$ is less than the smallest nonzero eigenvalue of ${L_{W^{*}\left(A^{\oplus n+1}\right)W}}(X)(X)$ guarantees that inequality (3.6) holds, hence inequality (3.5) holds.
Now for each $W\in{\mathcal{I}}_{B}(n+1)$, set
$\begin{array}[]{rllcl}\tilde{c}_{W}:=&\underset{c_{W}\in\mathbb{R}}{\sup}\ \ \ \ c_{W}\\ \mathrm{s.t.}&{L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&c_{W}\beta\\ c_{W}\beta^{*}&0\end{pmatrix}\succeq 0.\end{array}$
Note that our previous argument shows that $\tilde{c}_{W}>0$ for all $W\in{\mathcal{I}}_{B}(n+1)$. Additionally, again considering the Schur complement shows that for each fixed $W$ one has
${L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&\alpha\beta\\ \alpha\beta^{*}&0\end{pmatrix}\succeq 0$
for all $\alpha\in[0,\tilde{c}_{W}]$.
Now set
$\tilde{c}=\inf_{W\in{\mathcal{I}}_{B}(n+1)}\tilde{c}_{W}.$
If $\tilde{c}=\infty$, then we have that
${L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&\beta\\ \beta^{*}&0\end{pmatrix}\succeq 0$
for all $W\in{\mathcal{I}}_{B}(n+1)$, in which case the result follows. On the other hand, if $\tilde{c}<\infty$, then since ${\mathcal{I}}_{B}(n+1,r)$ is compact for each $r\leq n$ and since ${\mathcal{I}}_{B}(n+1)=\cup_{r\leq n+1}{\mathcal{I}}_{B}(n+1,r)$, a straightforward argument shows that there is some $W^{\prime}\in{\mathcal{I}}_{B}(n+1)$ such that
$\tilde{c}=\tilde{c}_{W^{\prime}}\qquad\mathrm{hence}\qquad\tilde{c}>0.$
Since $0<\tilde{c}\leq\tilde{c}_{W}$ for all $W\in{\mathcal{I}}_{B}(n+1)$ we conclude that
${L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&\tilde{c}\beta\\ \tilde{c}\beta^{*}&0\end{pmatrix}\succeq 0$
for all $W\in{\mathcal{I}}_{B}(n+1)$. Using Theorem 2.1 then shows that
$\begin{pmatrix}X&\tilde{c}\beta\\ \tilde{c}\beta^{*}&0\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}},$
as claimed. The reverse direction is a consequence of Item (1).
Item (3) follows from Items (1) and (2). ∎
Remark 3.2.
As in [24], the ability to take $\gamma=0\in\mathbb{R}^{g}$ in Lemma 3.1 (2) helps simplify the NC $*{}^{*}$ calculations used in the upcoming proof of Theorem 3.7.
We will soon define maximal $1$-dilations for a given element of a free spectrahedrop. Before doing so, we give two lemmas which together imply this upcoming notion is well-defined.
Lemma 3.3.
Let $A\in{SM_{d}(\mathbb{R})^{g}}$ and $B\in{SM_{d}(\mathbb{R})^{h}}$ and assume ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded free spectrahedrop. Given a tuples $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$ and $\beta\in M_{n\times 1}(\mathbb{R})^{g}$, the optimization problem
(3.7) $\begin{array}[]{rllcl}\tilde{\alpha}:=&\underset{\alpha\in\mathbb{R},\gamma\in\mathbb{R}^{g}}{\mathrm{\sup}}\ \ \ \ \alpha\\ \mathrm{s.t.}&{L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&\alpha\beta\\ \alpha\beta^{*}&\gamma\end{pmatrix}\succeq 0\\ \\ \mathrm{for\ all\ }&W\in{\mathcal{I}}_{B}(n+1)\end{array}$
achieves its maximum.
Proof.
Straightforward from the assumption that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is compact. ∎
Before presenting our next lemma, we introduce some notation. Given a matrix convex set $K$ and tuples $X\in SM_{n}(\mathbb{R})^{g}$ and $\beta\in M_{n\times 1}(\mathbb{R})^{g}$, define
(3.8) $\Gamma_{X,\beta}(K):=\left\{\gamma\in\mathbb{R}^{g}:\ \begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K\right\}.$
Lemma 3.4.
Let $K$ be a closed bounded matrix convex set. Fix $X\in SM_{n}(\mathbb{R})^{g}$ and $\beta\in M_{n\times 1}(\mathbb{R})^{g}$. Then the set $\Gamma_{X,\beta}(K)$ is a closed bounded convex set.
Proof.
If $\Gamma_{X,\beta}(K)$ is empty, then the result trivially holds. When $\Gamma_{X,\beta}(K)$ is nonempty, the fact that $\Gamma_{X,\beta}(K)$ is closed and bounded is immediate from the fact that $K$ is closed and bounded.
We now show that $\Gamma_{X,\beta}(K)$ is convex. To this end, let $\{\gamma_{1},\dots,\gamma_{k}\}\subseteq\Gamma_{X,\beta}(K)$ and let $\{c_{1},\dots,c_{k}\}$ be nonnegative constants such that $\sum_{i=1}^{k}c_{i}=1$. For each $i=1\dots,k$ set $V_{i}=\sqrt{c_{i}}I_{n+1}.$ Then we have
$\begin{pmatrix}X&\beta\\ \beta^{*}&\sum_{i=1}^{k}c_{i}\gamma^{i}\end{pmatrix}=\sum_{i=1}^{k}V_{i}^{*}\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma^{i}\end{pmatrix}V_{i}\qquad\mathrm{and}\qquad\sum_{i=1}^{k}V_{i}^{*}V_{i}=1.$
From the definition of $\Gamma_{X,\beta}(K)$, the above is a matrix convex combination of elements of $K$. Since $K$ is matrix convex, it follows that
$\begin{pmatrix}X&\beta\\ \beta^{*}&\sum_{i=1}^{k}c_{i}\gamma^{i}\end{pmatrix}\in K,$
from which we obtain $\sum_{i=1}^{k}c_{i}\gamma^{i}\in\Gamma_{X,\beta}(K)$. That is, $\Gamma_{X,\beta}$ is convex. ∎
3.1. Maximal 1-dilations in free spectrahedrops
We now present our definition of maximal $1$-dilations for elements of free spectradrops. We mention that our definition is a generalization of the definition of maximal $1$-dilations given in [24]. The definition used in [24] was itself inspired by works such as [20], [5], and [13].
Definition 3.5.
Given a closed bounded free spectrahedrop ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ and a tuple $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$, say the dilation
[…]
is a maximal $1$-dilation of $X$ if ${\hat{\beta}}\in M_{n\times 1}(\mathbb{R})^{g}$ is nonzero and the following two conditions hold:
  1. (1)
    The real number $1$ satisfies
    $\begin{array}[]{rllcl}1=&\underset{\alpha\in\mathbb{R},\gamma\in\mathbb{R}^{g}}{\max}\ \ \ \ \alpha\\ \mathrm{s.t.}&{L_{W^{*}\left(A^{\oplus n+1}\right)W}}\begin{pmatrix}X&\alpha{\hat{\beta}}\\ \alpha{\hat{\beta}}^{*}&\gamma\end{pmatrix}\succeq 0\\ \mathrm{for\ all}&W\in{\mathcal{I}}_{B}(n+1)\end{array}$
  2. (2)
    ${\hat{\gamma}}$ is an extreme point of the closed bounded convex set ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$.
Proposition 3.6.
Let ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ be a closed bounded free spectrahedrop and let $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$. If $X$ is not an Arveson extreme point of ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$, then there exists a
${\hat{Y}}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$
such that ${\hat{Y}}$ is a maximal $1$-dilation of $X$.
Proof.
The existence of maximal $1$-dilations follows from Lemma 3.1 together with Lemmas 3.3 and 3.4. In particular, Lemma 3.1 shows that if $X$ is not Arveson extreme, then there exists some nonzero $\beta\in{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$. Furthermore, Lemmas 3.1 and 3.3 together show that if one takes $\tilde{\alpha}$ to as defined in equation (3.7) and sets $\hat{\beta}=\tilde{\alpha}\beta$, then $\hat{\beta}$ is nonzero and Condition 4.5 (1) is satisfied. Combining this with Lemma 3.4 then shows that then ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ is a closed bounded convex set. Furthermore, Lemma 3.1 shows that ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ is nonempty. Thus, if one takes $\hat{\gamma}$ to be an extreme point of ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ then Condition 4.5 (2) is satisfied. It follows that
$\hat{Y}=\begin{pmatrix}X&{\hat{\beta}}\\ {\hat{\beta}}^{*}&{\hat{\gamma}}\end{pmatrix},$
is a maximal $1$-dilation of $X$. ∎
Our next theorem shows that maximal $1$-dilations of an element of a bounded free spectrahedrop reduce the dimension of the dilation subspace.
Theorem 3.7.
Let $A\in{SM_{d}(\mathbb{R})^{g}}$ and $B\in{SM_{d}(\mathbb{R})^{h}}$ such that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded real free spectrahedrop and let $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$. Assume $X$ is not an Arveson extreme point of ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$. Then there exists a maximal $1$-dilation ${\hat{Y}}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1)$ of $X$. Furthermore, any such ${\hat{Y}}$ satisfies
$\dim{\mathfrak{D}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{\mathfrak{D}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
Proof.
Before proceeding, we mention that the proof of this result follows the same flow as the proof of [24, Theorem 2.4]. Modifications are made to handle the fact that the kernel containment appearing in equation (3.2) must hold for all $W\in{\mathcal{I}}_{B}(n+1)$ for a tuple $\beta$ to be an element of ${{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$.
The existence of a maximal $1$-dilation of $X$ is proved in Proposition 3.6. Now, let
${\hat{Y}}=\begin{pmatrix}X&{\hat{\beta}}\\ {\hat{\beta}}^{*}&{\hat{\gamma}}\end{pmatrix}$
be a maximal $1$-dilation of $X$. Using Lemma 3.1 (3), it is sufficient to show that
$\dim{{\mathfrak{K}}_{\hat{Y}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
First consider the subspace
${{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}:=\left\{\eta\in M_{n\times 1}(\mathbb{R})^{g}\ :\ \exists\ \sigma\in\mathbb{R}^{g}\mathrm{\ s.t.\ }\begin{pmatrix}\eta\\ \sigma\end{pmatrix}\in{{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\right\}.$
In other words ${{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ is the projection $\iota$ of ${{\mathfrak{K}}_{Y}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$ defined by
${{\mathfrak{E}}_{Y}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}:=\iota({{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})})\mathrm{\ where\ }\iota\begin{pmatrix}\eta\\ \sigma\end{pmatrix}=\eta$
for $\eta\in M_{n\times 1}(\mathbb{R})^{g}$ and $\sigma\in\mathbb{R}^{g}$. We will show
$\dim{{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
If $\eta\in{{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$, then there exists a tuple $\tilde{\sigma}\in\mathbb{R}^{g}$ such that
$\begin{pmatrix}\eta^{*}&\tilde{\sigma}\end{pmatrix}\in{{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
From Lemma 3.1 (2), it follows that there is a real number $c>0$ so that setting $\sigma=c\tilde{\sigma}$ gives
$\begin{pmatrix}X&{\hat{\beta}}&c\eta\\ {\hat{\beta}}^{*}&{\hat{\gamma}}&\sigma\\ c\eta^{*}&\sigma^{*}&0\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}.$
Since ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is matrix convex it follows that
$\begin{pmatrix}1&0&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}X&{\hat{\beta}}&c\eta\\ {\hat{\beta}}^{*}&{\hat{\gamma}}&\sigma\\ c\eta^{*}&\sigma^{*}&0\end{pmatrix}\begin{pmatrix}1&0\\ 0&0\\ 0&1\end{pmatrix}=\begin{pmatrix}X&c\eta\\ c\eta^{*}&0\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}},$
so Lemma 3.1 (1) shows $\eta\in{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$. In particular this shows
(3.9) ${{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\subseteq{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})},$
hence
(3.10) $\dim{{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\leq\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
Now, assume towards a contradiction that
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\geq\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
We next show that this implies that there is a real number $c$ and a tuple $\sigma\in\mathbb{R}^{g}$ so that
(3.11) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}\begin{pmatrix}X&\hat{\beta}&c\hat{\beta}\\ \hat{\beta}^{*}&\hat{\gamma}&\sigma\\ c\hat{\beta}^{*}&\sigma&0\end{pmatrix}\succeq 0\qquad\mathrm{for\ all\ }W\in{\mathcal{I}}_{B}(n+2).$
and such that either $c\neq 0$ or $\sigma\neq 0$. To see this, observe that equation (3.9) implies that if $\dim{{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}=\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$, then we must have
${{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}={{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
In this case, using Lemma 3.1 (2) together with Theorem 2.1 (2) shows that there is a nonzero $c\in\mathbb{R}$ and some (possibly zero) $\sigma\in\mathbb{R}^{g}$ such that inequality (3.11) holds. On the other hand, if $\dim{{\mathfrak{E}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$, then there must exist tuples $\eta\in M_{n\times 1}(\mathbb{R})^{g}$ and $\sigma^{1},\sigma^{2}\in\mathbb{R}^{g}$ such that $\sigma^{1}\neq\sigma^{2}$ and so
[…]
In this case, setting $\sigma=\alpha(\sigma^{1}-\sigma^{2})$ for an appropriately chosen constant $\alpha\in\mathbb{R}$ and again using using Lemma 3.1 (2) together with Theorem 2.1 (2) shows that there is a (possibly zero) real number $c$ and a nonzero tuple $\sigma$ such that inequality (3.11) holds. Thus we have proved our claim that there is either a nonzero $c\in\mathbb{R}$ or a nonzero $\sigma\in\mathbb{R}^{g}$ such that inequality (3.11) holds.
We now use inequality (3.11) together with the NC $*{}^{*}$-decomposition to show that ${\hat{Y}}$ cannot be a maximal $1-$dilation of $X$, a contradiction to our definition of ${\hat{Y}}$. Applying the NC $*{}^{*}$-decomposition (up to canonical shuffles) shows that, for each fixed $W\in{\mathcal{I}}_{B}(n+2)$, inequality (3.11) holds if and only if ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(X)\succeq 0$ and the Schur complements
(3.13) $I_{n+2}-c^{2}Q_{W}\succeq 0,$
and
(3.14) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}({\hat{\gamma}})-Q_{W}$
-- $\left({\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\sigma)-cQ_{W}\right)^{*}\left(I-c^{2}Q_{W}\right)^{\dagger}\left({\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\sigma)-cQ_{W}\right)\succeq 0,$
where
(3.15) $Q_{W}:={\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}({\hat{\beta}}^{*}){L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(X)^{\dagger}{\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}({\hat{\beta}}).$
It follows that
(3.16) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\hat{\gamma})-Q_{W}\succeq 0,$
and
(3.17) $\ker[{L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\hat{\gamma})-Q_{W}]\subseteq\ker[{\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\sigma)-cQ_{W}].$
Inequalities (3.16) and (3.17) imply that for each fixed $W\in{\mathcal{I}}_{B}(n+2)$ there exists a real number $\alpha_{W}>0$ such that $0<\alpha\leq\alpha_{W}$ implies
(3.18) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\hat{\gamma})-Q_{W}\ \pm\ \alpha\left({\Lambda_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\sigma)-cQ_{W}\right)\succeq 0.$
Our next goal is to show that there is some real number $\tilde{\alpha}>0$ which is independent of $W\in{\mathcal{I}}_{B}(n+2)$ such that equation (3.18) holds for all $0<\alpha\leq\tilde{\alpha}$ and all $W\in{\mathcal{I}}_{B}(n+2)$. We note that the proof of this fact is similar to the proof of Lemma 3.1 (2), however, we include the details for the sake of completeness.
For each fixed $W$, set $\tilde{\alpha}_{W}$ to be the largest real number such that equation (3.18) holds for this fixed $W$ and for all $0<\alpha\leq\tilde{\alpha}_{W}$. Additionally define
(3.19) $\tilde{\alpha}=\min_{r\leq n+2}\ \ \inf_{W\in{\mathcal{I}}_{B}(n+2,r)}\ \ \tilde{\alpha}_{W}.$
A routine compactness argument shows that for each fixed $r\in\{1,2,\dots,n+2\}$ there is some $\tilde{W}_{r}\in{\mathcal{I}}_{B}(n+2,r)$ such that
$\tilde{\alpha}_{\tilde{W}_{r}}=\inf_{W\in{\mathcal{I}}_{B}(n+2,r)}\ \ \tilde{\alpha}_{W}.$
From this we can conclude that there is some some $\tilde{r}\in\{1,2,\dots,n+2\}$ such that $\tilde{\alpha}=\tilde{\alpha}_{\tilde{W}_{\tilde{r}}}.$ Recall from the previous part of the proof that $\tilde{\alpha}_{\tilde{W}_{\tilde{r}}}>0$. We conclude that if one sets $\tilde{\alpha}$ as in equation (3.19), then one indeed has that $\tilde{\alpha}>0$ and that that equation (3.18) holds for all $0<\alpha\leq\tilde{\alpha}$ and all $W\in{\mathcal{I}}_{B}(n+2)$.
After simplifying equation (3.18), it follows from the above that
(3.20) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}(\hat{\gamma}\pm\tilde{\alpha}\sigma)-(1\pm c\tilde{\alpha})Q_{W}\succeq 0$
for all $W\in{\mathcal{I}}_{B}(n+2)$. Furthermore, since $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$, using Theorem 2.1 (3) shows ${L_{W^{*}\left(A^{\oplus n+2}\right)W}}(X)\succeq 0$ for all $W\in{\mathcal{I}}_{B}(n+2)$. Combining this with equation (3.20) and another application of the Schur compliment shows that
(3.21) ${L_{W^{*}\left(A^{\oplus d(n+2)}\right)W}}\begin{pmatrix}X&\sqrt{1\pm c\tilde{\alpha}}\hat{\beta}\\ \sqrt{1\pm c\tilde{\alpha}}\hat{\beta}^{*}&\hat{\gamma}\pm\tilde{\alpha}\sigma\end{pmatrix}\succeq 0\qquad\mathrm{for\ all}\qquad W\in{\mathcal{I}}_{B}(n+2).$
[…]
(3.22) ${L_{W^{*}\left(A^{\oplus d(n+1)}\right)W}}\begin{pmatrix}X&\sqrt{1\pm c\tilde{\alpha}}\hat{\beta}\\ \sqrt{1\pm c\tilde{\alpha}}\hat{\beta}^{*}&\hat{\gamma}\pm\alpha\sigma\end{pmatrix}\succeq 0\qquad\mathrm{for\ all}\qquad W\in{\mathcal{I}}_{B}(n+1).$
It follows from Theorem 2.1 that
$\begin{pmatrix}X&\sqrt{1\pm c\tilde{\alpha}}\hat{\beta}\\ \sqrt{1\pm c\tilde{\alpha}}\hat{\beta}^{*}&\hat{\gamma}\pm\alpha\sigma\end{pmatrix}\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n+1).$
Recalling that ${\hat{Y}}$ was chosen to be a maximal $1$-dilation of $X$, we must have
$\sqrt{1\pm c\tilde{\alpha}}\leq 1$
It follows that $c\tilde{\alpha}=0$. Moreover, since $\tilde{\alpha}>0$, we must have $c=0$. From this we find
${\hat{\gamma}}\pm\tilde{\alpha}\sigma\in\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}).$
From our construction we have that if $c=0$, then $\sigma\neq 0$, so since $\tilde{\alpha}>0$, the above implies that ${\hat{\gamma}}$ is not an extreme point of the closed bounded convex set ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$. However, this is a contradiction to ${\hat{Y}}$ being a maximal $1$-dilation of $X$, since, from the definition of a maximal $1$-dilation, ${\hat{\gamma}}$ is an extreme point of ${\Gamma_{X,{\hat{\beta}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}$. Thus, the assumption that
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}\geq\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
leads to a contradiction. We conclude that
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{{\mathfrak{K}}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
Using Lemma 3.1 (3) then shows that
$\dim{\mathfrak{D}_{{\hat{Y}}}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}<\dim{\mathfrak{D}_{X}({{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}})}.$
That is, maximal $1$-dilations reduce the dimension of the dilation subspace, as claimed. ∎
4. Arveson extreme points and generalized free spectrahedra
We now handle the case of real generalized free spectrahedra. In this section, $\mathbb{F}=\mathbb{R}$ and ${\mathcal{H}}$ is a real separable infinite dimensional Hilbert space and $A\in SA({\mathcal{H}})^{g}$ is a tuple of compact self-adjoint operators on ${\mathcal{H}}$. A simple observation is that with these assumptions, if $X\in{\mathcal{D}}_{A}$ then $L_{A}(X)$ has a smallest nonzero eigenvalue. Using this fact allows for us to define maximal $1$-dilations for generalized free spectrahedra and to show that these maximal $1$-dilations in generalized free spectrahedra reduce the dimension of the dilation subspace. Though the argument in this setting is again similar to arguments used in [24] (and in this article in Section 3), details are given for the sake of completeness.
Lemma 4.1.
Let $A$ be a $g$-tuple of bounded self-adjoint compact operators on ${\mathcal{H}}$ and let $X\in SM_{n}(\mathbb{R})^{g}$ be a $g$-tuple of self-adjoint $n\times n$ matrices. Then
$\Lambda_{A}(X):=A_{1}\otimes X_{1}+\dots+A_{g}\otimes X_{g}$
is a compact operator self-adjoint on ${\mathcal{H}}\otimes\mathbb{R}^{n}$
Proof.
Straightforward. ∎
Lemma 4.2.
Let $Q\in B({\mathcal{H}})$ be a compact self-adjoint operator. Then $I-Q$ has a smallest nonzero eigenvalue.
Proof.
Since $Q$ is compact and self-adjoint, $Q$ is diagonalizable and only can have zero as a limit point of its spectrum. It follows that
$L_{A}(X):=I-\Lambda_{A}(X)$
is diagonalizable and can only have one as a limit point of its spectrum. Therefore, $I-Q$ has a smallest nonzero eigenvalue. ∎
Corollary 4.3.
Let $A$ be a $g$-tuple of compact self-adjoint operators on ${\mathcal{H}}$ such that ${\mathcal{D}}_{A}$ is a generalized free spectrahedron and let $X\in SM_{n}(\mathbb{R})^{g}$ be any $g$-tuple of self-adjoint $n\times n$ matrices. Then $L_{A}(X)$ has a smallest nonzero eigenvalue.
As an immediate consequence $L_{A}(X)^{\dagger}$ is a bounded self-adjoint operator on ${\mathcal{H}}\otimes\mathbb{R}^{n}$.
4.1. The dilation subspace of generalized free spectrahedra
Similar to as was done for free spectrahedrops, we now describe an alternative characterization of the dilation subspace for generalized free spectrahedra. To this end, define the subspace ${{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}$ by
${{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}=\{\beta\in M_{n\times 1}(\mathbb{R})^{g}|\ \ker L_{A}(X)\subseteq\ker\Lambda_{A}(\beta^{*})\}.$
Lemma 4.4.
Let ${\mathcal{D}}_{A}$ be a generalized free spectrahedron and let $X\in{\mathcal{D}}_{A}(n)$.
  1. (1)
    If $\beta\in M_{n\times 1}(\mathbb{R})^{g}$ and
    $Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in{\mathcal{D}}_{A}(n+1)$
    is a $1$-dilation of $X$, then $\beta\in{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}$.
  2. (2)
    Let $\beta\in M_{n\times 1}(\mathbb{R})^{g}$. Then $\beta\in{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}$ if and only if there is a tuple $\gamma\in{\mathcal{D}}_{A}(1)$ real number $c_{\gamma}>0$ such that
    $\begin{pmatrix}X&c_{\gamma}\beta\\ c_{\gamma}\beta^{*}&\gamma\end{pmatrix}\in{\mathcal{D}}_{A}(n+1).$
    In particular, one may take $\gamma=0\in\mathbb{R}^{g}$.
  3. (3)
    One has
    ${\mathfrak{D}_{X}({\mathcal{D}}_{A})}={{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}.$
    As a consequence, $X$ is an Arveson extreme point of ${\mathcal{D}}_{A}$ if and only if
    $\dim{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}=0.$
Proof.
Corollary 4.3 shows that $L_{A}(\gamma)^{\dagger}$ is a well-defined bounded operator on ${\mathcal{H}}$. From this point, the proof if Item (1) is essentially identical to the proof of Lemma 3.1 (1).
We now prove Item (2). Note that $L_{A}(0)=I_{\mathcal{H}}$, so using the Schur complement shows
$Y_{0}=\begin{pmatrix}X&c\beta\\ c\beta^{*}&0\end{pmatrix}\in{\mathcal{D}}_{A}(n+1),$
if and only if
(4.1) $L_{A}(X)-c^{2}\Lambda_{A}(\beta)\Lambda_{A}(\beta^{*})\succeq 0.$
Using Corollary 4.3 shows that $L_{A}(X)$ has a smallest nonzero eigenvalue, so we may pick $c$ small enough so that $\|c^{2}\Lambda_{A}(\beta)\Lambda_{A}(\beta^{*})\|_{2}$ is less than the smallest nonzero eigenvalue of $L_{A}(X)$. Furthermore, $\beta\in{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}$ implies $\ker L_{A}(X)\subseteq\ker\Lambda_{A}(\beta)\Lambda_{A}(\beta^{*})$. Thus this choice of $c$ then guarantees that inequality (4.1) holds, hence $Y_{0}\in{\mathcal{D}}_{A}(n+1)$. The reverse direction is a consequence of Item (1).
Item (3) follows from Items (1) and (2). ∎
4.2. Maximal 1-dilations for generalized free spectrahedra.
We now present our notion of maximal $1$-dilations for generalized free spectrahedra.
Definition 4.5.
Given a bounded generalized free spectrahedron ${\mathcal{D}}_{A}$ and a tuple $X\in{\mathcal{D}}_{A}(n)$, say the dilation
$\hat{Y}=\begin{pmatrix}X&{\hat{\beta}}\\ {\hat{\beta}}^{*}&{\hat{\gamma}}\end{pmatrix}\in{\mathcal{D}}_{A}(n+1)$
is a maximal $1$-dilation of $X$ if ${\hat{\beta}}\in M_{n\times 1}(\mathbb{R})^{g}$ is nonzero and the following two conditions hold:
  1. (1)
    The real number $1$ satisfies
    $\begin{array}[]{rllcl}1=&\underset{\alpha\in\mathbb{R},\gamma\in\mathbb{R}^{g}}{\max}\ \ \ \ \alpha\\ \mathrm{s.t.}&L_{A}\begin{pmatrix}X&\alpha{\hat{\beta}}\\ \alpha{\hat{\beta}}^{*}&\gamma\end{pmatrix}\succeq 0\end{array}$
  2. (2)
    ${\hat{\gamma}}$ is an extreme point of the closed bounded convex set ${\Gamma_{A,{\hat{\beta}}}({\mathcal{D}}_{A})}$ where ${\Gamma_{A,{\hat{\beta}}}({\mathcal{D}}_{A})}$ is as defined in equation (3.8).
We now show that maximal $1$-dilations in generalized free spectrahedra reduce the dimension of the dilation subspace.
Theorem 4.6.
Let $A\in B({\mathcal{H}})^{g}$ be a $g$-tuple of compact self-adjoint operators on ${\mathcal{H}}$ such that ${\mathcal{D}}_{A}$ is a bounded real free spectrahedron and let $X\in{\mathcal{D}}_{A}(n)$. Assume $X$ is not an Arveson extreme point of ${\mathcal{D}}_{A}$. Then there exists a nontrivial maximal 1-dilation ${\hat{Y}}\in{\mathcal{D}}_{A}(n+1)$ of $X$. Furthermore, any such ${\hat{Y}}$ satisfies
$\dim{\mathfrak{D}_{{\hat{Y}}}({\mathcal{D}}_{A})}<\dim{\mathfrak{D}_{X}({\mathcal{D}}_{A})}.$
Proof.
The existence of maximal $1$-dilations in a bounded generalized free spectrahedron follows from a routine compactness argument together with Lemma 3.4.
Now, let ${\hat{Y}}$ be a maximal $1$-dilation of $X$. Using Lemma 4.4 (3), it is sufficient to show
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({\mathcal{D}}_{A})}<\dim{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}.$
Assume towards a contradiction that
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({\mathcal{D}}_{A})}\geq\dim{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})}.$
Following the same argument as was given at the beginning of the proof of Theorem 3.7 shows that there exists a real number $c$ and a tuple $\sigma\in\mathbb{R}^{g}$ so that
(4.2) $L_{A}\begin{pmatrix}X&\hat{\beta}&c\hat{\beta}\\ \hat{\beta}^{*}&\hat{\gamma}&\sigma\\ c\hat{\beta}^{*}&\sigma&0\end{pmatrix}\succeq 0.$
and such that either $c\neq 0$ or $\sigma\neq 0$.
Now, applying the NC $*{}^{*}$-decomposition (up to canonical shuffles) shows that inequality (4.2) holds if and only if $L_{A}(X)\succeq 0$ and the Schur complements
(4.3) $I_{\mathcal{H}}-c^{2}Q\succeq 0,$
and
(4.4) $L_{A}({\hat{\gamma}})-Q-\left(\Lambda_{A}(\sigma)-cQ\right)^{*}\left(I_{\mathcal{H}}-c^{2}Q\right)^{\dagger}\left(\Lambda_{A}(\sigma)-cQ\right)\succeq 0,$
where
(4.5) $Q:=\Lambda_{A}({\hat{\beta}}^{*})L_{A}(X)^{\dagger}\Lambda_{A}({\hat{\beta}}).$
It follows that
(4.6) $L_{A}(\hat{\gamma})-Q\succeq 0,$
and
(4.7) $\ker[L_{A}(\hat{\gamma})-Q]\subseteq\ker[\Lambda_{A}(\sigma)-cQ].$
Recall from Corollary 4.3 and Lemma 4.1 that $L_{A}(X)^{\dagger}$ is a bounded self-adjoint operator and $\Lambda_{A}({\hat{\beta}})$ and $\Lambda_{A}({\hat{\beta}}^{*})$ are compact operators. It follows that $Q$ is a compact self-adjoint operator. Therefore
$L_{A}(\hat{\gamma})-Q=I_{\mathcal{H}}-(\Lambda_{A}(\hat{\gamma})+Q),$
is the identity minus a compact self-adjoint operator and by Lemma 4.2 has a smallest nonzero eigenvalue. Therefore, picking $\tilde{\alpha}>0$ so that $\tilde{\alpha}\|\Lambda_{A}(\sigma)-cQ\|$ is smaller than the smallest nonzero eigenvalue of $L_{A}(\hat{\gamma})-Q$ and using inequalities (4.6) and (4.7) guarantees
$L_{A}(\hat{\gamma})-Q\ \pm\ \tilde{\alpha}\left(\Lambda_{A}(\sigma)-cQ\right)\succeq 0.$
It follows from the above that
(4.8) $\begin{array}[]{rllcl}&L_{A}(\hat{\gamma}\pm\tilde{\alpha}\sigma)-(1\pm c\tilde{\alpha})Q\\ =&L_{A}(\hat{\gamma}\pm\alpha\sigma)-\left(\Lambda_{A}(\sqrt{1\pm c\tilde{\alpha}}\hat{\beta}^{*})L_{A}(X)^{\dagger}\Lambda_{A}(\sqrt{1\pm c\tilde{\alpha}}\hat{\beta})\right)&\succeq&0.\end{array}$
Since $L_{A}(X)\succeq 0$, equation (4.8) implies
(4.9) $L_{A}\begin{pmatrix}X&\sqrt{1\pm c\tilde{\alpha}}\hat{\beta}\\ \sqrt{1\pm c\tilde{\alpha}}\hat{\beta}^{*}&\hat{\gamma}\pm\tilde{\alpha}\sigma\end{pmatrix}\succeq 0.$
Recalling that ${\hat{Y}}$ is a maximal $1$-dilation of $X$, we must have
$\sqrt{1\pm c\tilde{\alpha}}\leq 1.$
It follows that $c\tilde{\alpha}=0$. Since $\tilde{\alpha}>0$, it follows that $c=0$. From our construction, this in turn implies that $\sigma\neq 0$. But then equation (4.9) implies that
$\hat{\gamma}\pm\tilde{\alpha}\sigma\in{\Gamma_{X,{\hat{\beta}}}({\mathcal{D}}_{A})},$
which contradicts the fact that $\hat{\gamma}$ is an extreme point of the convex set ${\Gamma_{X,{\hat{\beta}}}({\mathcal{D}}_{A})}$. We conclude that
$\dim{{\mathfrak{K}}_{{\hat{Y}}}({\mathcal{D}}_{A})}<\dim{{\mathfrak{K}}_{X}({\mathcal{D}}_{A})},$
from which we can use Lemma 4.4 (3) to conclude that
${\mathfrak{D}_{{\hat{Y}}}({\mathcal{D}}_{A})}<{\mathfrak{D}_{X}({\mathcal{D}}_{A})},$
as claimed. ∎
5. From reals to complexes
We now show how to reduce spanning over the complexes to spanning over the reals. A major goal of the section is to generalize [24, Section 3] which considers free spectrahedra to general matrix convex sets. We note that while [24, Section 3] restricted to free spectrahedra, many of the proofs there only relied on matrix convexity and immediately generalize to our present setting. Proofs of such results will be given in Appendix 8.
A tool we use in throughout the section is the standard embedding of a tuple of complex self-adjoint matrices to a tuple of real symmetric matrices. To this end, given an natural number $m$, we define $U_{m}\in\mathbb{C}^{2m\times 2m}$ to be the unitary
(5.1) $U_{m}:=\frac{\sqrt{2}}{2}\begin{pmatrix}I_{m}&-iI_{m}\\ -iI_{m}&I_{m}\end{pmatrix}\in M_{2m}(\mathbb{C}).$
A key fact is that if $X\in SM_{m}(\mathbb{C})^{g}$ with real and imaginary parts $X_{\mathbb{R}}$ and $X_{\mathbb{C}}$ then
$U_{m}^{*}(X\oplus\overline{X})U_{m}=\begin{pmatrix}X_{\mathbb{R}}&X_{\mathbb{C}}\\ X_{\mathbb{C}}^{T}&X_{\mathbb{R}}\end{pmatrix}.$
We note that, since $X$ is self-adjoint, $X_{\mathbb{R}}$ is symmetric while $X_{\mathbb{C}}$ is skew symmetric.
We begin by relating the Arveson extreme points of a matrix convex set over the complexes to the Arveson extreme points of the associated matrix convex set over the reals.
Lemma 5.1.
Let $K\subseteq SM(\mathbb{C})^{g}$ be a matrix convex set such that $K=\overline{K}$, and let $K^{\mathbb{R}}$ denote $K\cap SM(\mathbb{R})^{g}$.
  1. (1)
    $X\in\partial^{\mathrm{Arv}}(K)$ if and only if $\overline{X}\in\partial^{\mathrm{Arv}}(K)$.
  2. (2)
    If $X\in K^{\mathbb{R}}$, then $X\in\partial^{\mathrm{Arv}}(K)$ if and only if $X\in\partial^{\mathrm{Arv}}(K^{\mathbb{R}})$.
Proof.
See Appendix 8.1. ∎
Remark 5.2.
It may not be the case that the real free extreme points of $K^{\mathbb{R}}$ are free extreme points of $K$. This is due to the issue of reducibility over $\mathbb{R}$ vs $\mathbb{C}$. For example, if $X\in K(n)$ is a free extreme point of $K$ with $X_{\mathbb{C}}\neq 0$, then the tuple $U_{n}(X\oplus\overline{X})U_{n}$ is reducible over $\mathbb{C}$ while it is irreducible over $\mathbb{R}$. Therefore, $U_{n}(X\oplus\overline{X})U_{n}$ is a free extreme point of $K^{\mathbb{R}}$, but not a free extreme point of $K$.
5.1. Proof of Theorem 2.6
Proof.
We first show that that $K=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K))$ implies $K^{\mathbb{R}}=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K^{\mathbb{R}}))$. To this end, let $X\in K^{\mathbb{R}}\subseteq K$. By assumption, there exists a tuple
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K(m)$
for some appropriate $m$ such that $Y\in\partial^{\mathrm{Arv}}(K)$. Since $K$ is closed under complex conjugation, we find that $X$ dilates to
$U_{m}^{*}(Y\oplus\overline{Y})U_{m}=\begin{pmatrix}Y_{\mathbb{R}}&Y_{\mathbb{C}}\\ Y_{\mathbb{C}}^{T}&Y_{\mathbb{R}}\end{pmatrix}=\begin{pmatrix}X&\beta_{\mathbb{R}}&0&\beta_{\mathbb{C}}\\ \beta_{\mathbb{R}}^{T}&\gamma_{\mathbb{R}}&-\beta_{\mathbb{C}}^{T}&\gamma_{\mathbb{C}}\\ 0&-\beta_{\mathbb{C}}&X&\beta_{\mathbb{R}}\\ \beta_{\mathbb{C}}^{T}&\gamma_{\mathbb{C}}^{T}&\beta_{\mathbb{R}}^{T}&\gamma_{\mathbb{R}}\end{pmatrix}\in K^{\mathbb{R}}(2m).$
Since the Arveson boundary is closed under unitary conjugation and since the above tuple is real valued, we can apply Lemma 5.1 (2) to conclude that
$U_{m}^{*}(Y\oplus\overline{Y})U_{m}\in\partial^{\mathrm{Arv}}(K^{\mathbb{R}})$
is an Arveson dilation of $X$ in $K^{\mathbb{R}}$. It follows that $K^{\mathbb{R}}=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K^{\mathbb{R}}))$.
We now show that $K^{\mathbb{R}}=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K^{\mathbb{R}}))$ implies $K=\mathrm{co}^{\mathrm{mat}}(\partial^{\mathrm{free}}(K))$. To this end let $X\in K(n)$. Since $K$ is closed under complex conjugation, we have $X\oplus\overline{X}$ in $K$, hence
$Z:=U_{n}^{*}\begin{pmatrix}X&0\\ 0&\overline{X}\end{pmatrix}U_{n}=\begin{pmatrix}X_{\mathbb{R}}&X_{\mathbb{C}}\\ X_{\mathbb{C}}^{T}&X_{\mathbb{R}}\end{pmatrix}\in{K^{\mathbb{R}}}.$
By assumption $K^{\mathbb{R}}$ is the matrix convex hull of its free extreme points, so $Z$ dilates to an Arveson extreme point $Y\in\partial^{\mathrm{Arv}}(K^{\mathbb{R}}(m))$ for some appropriate $m$. Again using Lemma 5.1 (2) shows that $Y\in\partial^{\mathrm{Arv}}(K)$. By decomposing $Y$ into its irreducible (over $\mathbb{C}$) components, we obtain a matrix convex combination of free extreme points of $K$ whose sum of sizes is at most $m$. ∎
5.2. The real part of a complex free spectrahedrop/generalized free spectrahedron
To use Theorem 2.6 to reduce to the real case in the proof of 2.4, we need to prove that $K^{\mathbb{R}}$ is a real free spectrahedrop/real generalized free spectrahedron. This is handled by the following lemma.
Lemma 5.3.
Let $K\subseteq SM(\mathbb{C})^{g}$ be a complex free spectrahedrop/generalized free spectrahedron. Then $K^{\mathbb{R}}=K\cap SM(\mathbb{R})^{g}$ is a real free spectrahedrop/generalized free spectrahedron.
Proof.
The result for free spectrahedrops follows from [24, Lemma 3.3] which has the consequence that the real part of a free spectrahedron closed under complex conjugation is indeed a real free spectrahedron. The proof for generalized free spectrahedra is essentially the same as the proof of [24, Lemma 3.3]. Details are given in Appendix 8.2. ∎
Finally, for the dimension bound on the sum of sizes of free extreme points required for a free Caratheodory expansion that is presented in Theorem 2.4, we extend [24, Lemma 3.2] from free spectrahedra to general matrix convex sets.
Lemma 5.4.
Let $K\subseteq SM(\mathbb{F})^{g}$ be a matrix convex set. Then we have the following:
  1. (1)
    Let $X\in K(n)$ and $Y\in K(m)$. Then
    ${\mathfrak{D}_{X\oplus Y}(K)}=\left\{\left(\beta^{*}\ \sigma^{*}\right)^{*}\in M_{(n+m)\times 1}(\mathbb{F})^{g}|\ \beta\in{\mathfrak{D}_{X}(K)}\mathrm{\ and\ }\sigma\in{\mathfrak{D}_{Y}(K)}\right\},$
    and
    $\dim{\mathfrak{D}_{X\oplus Y}(K)}=\dim{\mathfrak{D}_{X}(K)}+\dim{\mathfrak{D}_{Y}(K)}.$
  2. (2)
    Let $X\in K(n)$ and let $U\in M_{n}(\mathbb{F})$ be a unitary.
    ${\mathfrak{D}_{X}(K)}=U^{*}{\mathfrak{D}_{U^{*}XU}(K)}\qquad\mathrm{and}\qquad\dim{\mathfrak{D}_{X}(K)}=\dim{\mathfrak{D}_{U^{*}XU}(K)}.$
  3. (3)
    If $K$ is closed under complex conjugation, then for any $X\in K$,
    ${\mathfrak{D}_{X}(K)}=\overline{{\mathfrak{D}_{\overline{X}}(K)}}\qquad\mathrm{and}\qquad\dim{\mathfrak{D}_{X}(K)}=\dim{\mathfrak{D}_{\overline{X}}(K)}.$
  4. (4)
    If $K$ is closed under complex conjugation, then for any $X\in K(n)$,
    $\dim{\mathfrak{D}_{U_{n}^{*}\left(X\oplus\overline{X}\right)U_{n}}(K^{\mathbb{R}})}=2\dim{\mathfrak{D}_{X}(K)},$
    where $U_{n}$ is the unitary defined in equation (5.1).
Proof.
See Appendix 8.3. ∎
6. Free spectrahedrops versus generalized free spectrahedra
We now discuss a class of generalized free spectrahedra that are not free spectrahedrops. The class arises by considering the polar duals of the matrix convex sets introduced in [22] which have no free extreme points. Following the notation in [22], let ${\mathcal{H}}$ be a real separable infinite dimensional Hilbert space and let $A\in SA({\mathcal{H}})^{g}$. For each $n\in{\mathbb{N}}$ define the set $K_{A}(n)\subseteq SM_{n}(\mathbb{R})^{g}$ by
(6.1) $K_{A}(n)=\{Y\in SM_{n}(\mathbb{R})^{g}|Y=V^{*}(I_{\mathcal{H}}\otimes X)V\mathrm{\ for\ some\ isometry\ }V:\mathbb{R}^{n}\to\oplus_{1}^{\infty}{\mathcal{H}}\}.$
We then define $K_{A}\subseteq SM(\mathbb{R})^{g}$ by
(6.2) $K_{A}=\cup_{n}K_{A}(n).$
We call $K_{A}$ the noncommutative convex hull of X. The set $K_{A}$ is closely related to the matrix range of $A$, see [47] for further discussion. For a $g$-tuple $A\in SA({\mathcal{H}})^{g}$, say 00 is in the finite interior of $K_{A}$ if there exist an integer $d\in{\mathbb{N}}$ and a nonzero vector $v\in({\mathcal{H}})^{d}$ such that.
$v^{*}(I_{d}\otimes A)v=0\in\mathbb{R}^{g}.$
Theorem 6.1.
Let $A\in SA({\mathcal{H}})^{g}$ be a tuple of compact operators such that ${\mathcal{D}}_{A}$ is a bounded generalized free spectrahedron. Additionally assume that $A$ has no finite dimensional reducing subspaces and that 00 is in the finite interior of $K_{A}$. Then ${\mathcal{D}}_{A}$ is not a free spectrahedrop.
Proof.
Using [22] shows that $K_{A}$ is a closed bounded matrix convex set that has no free extreme points. It then follows from Theorem 2.4 that $K_{A}$ cannot be a free spectrahedrop. On the other hand, combining the discussion in [47, Example 4.6] with [12, Proposition 3.1] shows that $K_{A}^{\circ}={\mathcal{D}}_{A}$. Furthermore, using [35, Lemma 4.2] (also see [12, Lemma 3.2]) shows that since $0\in K_{A}$, we have that $(K_{A}^{\circ})^{\circ}=K_{A}$, hence ${\mathcal{D}}_{A}^{\circ}=K_{A}$. To complete the proof note that if ${\mathcal{D}}_{A}$ was a closed bounded free spectrahedrop, then [35, Theorem 4.11] would imply that ${\mathcal{D}}_{A}^{\circ}=K_{A}$ is also a closed bounded free spectrahedrop. We have already shown that $K_{A}$ is not a free spectrahedrop, so it follows that ${\mathcal{D}}_{A}$ cannot be a free spectraherop. ∎
As it turns out, the precise examples of tuples $A$ such that $A$ has no finite dimensional reducing subspaces and $K_{A}$ contains 00 in the finite interior considered in [22, Proposition 4.1] do not have the property that ${\mathcal{D}}_{A}$ is bounded. However, it is possible to lightly modify these example to make them appropriate for our current use. Similar to [22, Proposition 4.1], let ${\mathcal{H}}=\ell^{2}({\mathbb{N}})$ and let ${\mathcal{H}}_{2}=\ell^{2}(\{1,2\})$ be real Hilbert spaces. Additionally, given a weight vector $w=(w_{1},w_{2},\dots)\in\mathbb{R}^{\infty}$, define the weighted forward shift $S_{w}:{\mathcal{H}}\to{\mathcal{H}}$ by
$S_{w}v=(0,w_{1}v_{1},w_{2}v_{2},\dots)$
for all $v=(v_{1},v_{2},\dots)\in{\mathcal{H}}$.
Proposition 6.2.
Let $A_{1}=\mathrm{diag}(\lambda_{1},\lambda_{2},\dots)$ where the $\lambda_{i}$ are nonzero real numbers converging to 00 with distinct norms and where that $\lambda_{1}>0$ and that $\lambda_{2}<0$. Also let $S_{w}$ be a weighted shift where the weight vector $w\in\mathbb{R}^{\infty}$ is chosen so $w_{i}\neq 0$ for all $i$ and such that $S_{w}$ is compact. Set
$A_{2}=S_{w}+S_{w}^{*}.$
Then $(A_{1},A_{2})$ is a tuple of compact operators on ${\mathcal{H}}$ which has no finite dimensional reducing subspaces such that $K_{A}$ contains 00 in its finite interior and such that ${\mathcal{D}}_{A}$ is bounded.
Proof.
The proof that $A$ has no finite dimensional reducing subspaces is identical to proof appearing in [22, Proposition 5.1]. We next show that ${\mathcal{D}}_{A}$ is bounded. To this end, consider the inclusion map $\iota:{\mathcal{H}}_{2}\to{\mathcal{H}}$. By identifying ${\mathcal{H}}_{2}$ with $\mathbb{R}^{2}$ we have
$\iota^{*}A\iota=\left(\begin{pmatrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{pmatrix},\begin{pmatrix}0&w_{1}\\ w_{1}&0\end{pmatrix}\right)$
Since $\lambda_{1}>0$ and $\lambda_{2}<0$ and $w_{1}\neq 0$, no linear combination of $\iota^{*}A_{1}\iota$ and $\iota^{*}A_{2}\iota$ can be positive semidefinite, from which it follows that ${\mathcal{D}}_{\iota^{*}A\iota}$ is bounded. It is straightforward to check that
${\mathcal{D}}_{A}\subseteq{\mathcal{D}}_{\iota^{*}A\iota},$
so we conclude that ${\mathcal{D}}_{A}$ is bounded.
It remains to show that 00 is in the finite interior of $K_{A}$. Set
$v_{1}=\begin{pmatrix}\sqrt{-\lambda_{2}}\\ 0\end{pmatrix}{\quad\quad\mathrm{and}\quad\quad}v_{2}=\begin{pmatrix}0\\ \sqrt{\lambda_{1}}\end{pmatrix}$
Then one can verify that
$v_{1}^{*}(\iota^{*}A_{1}\iota)v_{1}+v_{2}^{*}(\iota^{*}A_{1}\iota)v_{2}=v_{1}^{*}(\iota^{*}A_{2}\iota)v_{1}+v_{2}^{*}(\iota^{*}A_{2}\iota)v_{2}=0.$
It follows that if one sets
$w=\begin{pmatrix}\iota v_{1}\\ \iota v_{2}\end{pmatrix}\in({\mathcal{H}})^{2},$
then $w^{*}(I_{2}\otimes A)w=0\in\mathbb{R}^{2}$. That is 00 is in the finite interior of $K_{A}$. ∎
6.1. Thoughts on “generalized” free spectrahedrops
To end the article, we briefly discuss the case of generalized free spectrahedrops. From our results, it is natural to wonder if it is necessary to consider finite dimensional defining tuples when working with free spectrahedrops. That is, suppose ${\mathcal{H}}$ is an separable infinite dimensional Hilbert space and suppose $A\in SA({\mathcal{H}})^{g}$ and $B\in SA({\mathcal{H}})^{h}$ are tuples of compact self-adjoint operators on ${\mathcal{H}}$ such that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded “generalized” free spectrahedrop that is closed under complex conjugation. One may wonder if ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is the matrix convex hull of its free extreme points.
It is unclear how to extend our approach to this setting. A key issue is that in the proof of [19, Theorem 3.2], given an element of $X\in{{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}(n)$ one considers a map
$W\to tr(W)I_{n}-\sum_{i=1}^{g}\langle A_{i},W\rangle X_{i}$
defined on $B({\mathcal{H}})$. Of course, such a map is not defined for general operators in $B({\mathcal{H}})$, which leads to difficulties.
Aside from this, we mention that other problems do appear to arise when one makes the jump from free spectrahedrops to generalized free spectrahedrops. For example, [35, Theorem 4.11] shows that if ${\mathcal{H}}$ is finite dimensional and ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a closed bounded free spectrahedrop, then ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is the projection of a bounded free spectrahedron. That is, one can assume that ${\mathcal{D}}_{(A,B)}$ is bounded if ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is closed and bounded. The proof of this result constructs a bounded free spectrahedron $K$ such that ${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}$ is a projection of $K$. However, the number of variables in $K$ depends on the dimension of ${\mathcal{H}}$. Thus, if one attempts to naively extend this proof to generalized free spectrahedrops, one would encounter dimension free sets in infinitely many variables.
Working with dimension free sets in infinitely many variables causes a number of challenges. For example, this would cause the approach of successive maximal $1$-dilations to fail. A critical aspect of this approach is that the dimension of the dilation subspace is bounded above by $ng$, hence is finite.
Appendix II contains proofs of several of the lemmas presented in Section 5
8.1. Proof of Lemma 5.1
Proof.
The proof follows the same approach as [24, Lemma 3.1]. We first prove item (1). Since $K$ is closed under complex conjugation, for any $\beta\in M_{n\times 1}(\mathbb{C})^{g}$ and $\gamma\in\mathbb{R}^{g}$, we find
$\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K\qquad\mathrm{if\ and\ only\ if}\qquad Y=\begin{pmatrix}\overline{X}&\overline{\beta}\\ \overline{\beta}^{*}&\overline{\gamma}\end{pmatrix}\in K.$
Therefore, there is a nontrivial dilation of $X$ in $K$ if and only if there is a nontrivial dilation of $\overline{X}$ in $K$.
We next prove item (2). Let $X\in K^{\mathbb{R}}$ and observe if $X$ cannot be trivially dilated over $\mathbb{C}$, then $X$ also cannot be trivially dilated over $\mathbb{R}$, hence it is clear that $X\in\partial^{\mathrm{Arv}}(K)$ implies $X\in\partial^{\mathrm{Arv}}(K^{\mathbb{R}})$. Now assume $X\in\partial^{\mathrm{Arv}}(K^{\mathbb{R}})$ and suppose $X$ dilates (over $\mathbb{C}$) to
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K(m).$
We will show that $\beta=0$.
Since $K$ is closed under complex conjugation, $X$ must also dilate to
$\overline{Y}=\begin{pmatrix}X&\overline{\beta}\\ \overline{\beta}^{*}&\overline{\gamma}\end{pmatrix}\in K(m),$
hence $X$ dilates to $Y\oplus\overline{Y}\in K(2m)$. Now, let $Y_{\mathbb{R}},\beta_{\mathbb{R}},\gamma_{R}$ and $Y_{\mathbb{C}},\beta_{\mathbb{C}},\gamma_{C}$ denote the real and imaginary parts of $Y,\beta,\gamma$, respectively. Then since $K$ is closed under unitary conjugation, we obtain
$U_{m}^{*}\begin{pmatrix}Y&0\\ 0&\overline{Y}\end{pmatrix}U_{m}=\begin{pmatrix}Y_{\mathbb{R}}&Y_{\mathbb{C}}\\ Y_{\mathbb{C}}^{T}&Y_{\mathbb{R}}\end{pmatrix}=\begin{pmatrix}X&\beta_{\mathbb{R}}&0&\beta_{\mathbb{C}}\\ \beta_{\mathbb{R}}^{T}&\gamma_{\mathbb{R}}&-\beta_{\mathbb{C}}^{T}&\gamma_{\mathbb{C}}\\ 0&-\beta_{\mathbb{C}}&X&\beta_{\mathbb{R}}\\ \beta_{\mathbb{C}}^{T}&\gamma_{\mathbb{C}}^{T}&\beta_{\mathbb{R}}^{T}&\gamma_{\mathbb{R}}\end{pmatrix}\in K^{\mathbb{R}}(2m).$
Here we have used the fact that $X$ is real valued and that $Y_{\mathbb{C}}$ is skew-symmetric, which follows from the fact that $Y$ is self-adjoint. Now, since $X$ is an Arveson extreme point of $K^{\mathbb{R}}$, we know that $X$ cannot be nontrivially dilated in $K^{\mathbb{R}}$, hence $\beta_{\mathbb{R}}=\beta_{\mathbb{C}}=0$. From this we obtain $\beta=0$ which proves that $X$ is an Arveson extreme point of $K$. ∎
8.2. Proof of Lemma 5.3
Proof.
First suppose $K={{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}\subseteq SM(\mathbb{C})^{g}$ is a complex free spectrahedrop where $A\in SM_{d}(\mathbb{C})^{g}$ and $B\in SM_{d}(\mathbb{C})^{h}$. Observe that
${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}\cap SM(\mathbb{R})^{g}=\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A\oplus\bar{A},B\oplus\bar{B})},$
as
$L_{(A,B)}(X,Y)\succeq 0\qquad\mathrm{if\ and\ only\ if}\qquad L_{\left(A\oplus\overline{A},B\oplus\overline{B}\right)}(X,Y)$
when $X,Y$ are real valued. It follows that
${{\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{(A,B)}}}\cap SM(\mathbb{R})^{g}=\mathrm{proj}_{x}\hskip 1.42271pt{\mathcal{D}}_{U_{d}^{*}\left(A\oplus\bar{A},B\oplus\bar{B}\right)U_{d}},$
where $U_{d}$ is the unitary defined in equation (5.1). Noting that the tuple $U_{d}^{*}\left(A\oplus\bar{A},B\oplus\bar{B}\right)U_{d}$ is real valued completes the proof.
Now let $K={\mathcal{D}}_{A}$ be a generalized free spectrahedron. Without loss of generality, we take the underlying complex separable infinite dimensional Hilbert space ${\mathcal{H}}$ to be $\ell^{2}(\mathbb{N})$. Then as above, since $K$ is closed under complex conjugation, for $X\in M_{n}(\mathbb{R})$, we have
$L_{A}(X)\succeq 0\qquad\mathrm{if\ and\ only\ if}\qquad L_{\overline{A}}(X)\succeq 0,$
where $\overline{A}=(\overline{A}_{1},\dots,\overline{A}_{g})\in SA(\ell^{2}(\mathbb{N}))^{g}$ and for each $i=1,\dots,g$ the operator $\overline{A}_{i}$ is defined by $\overline{A_{i}}v=\overline{A_{i}\overline{v}}$ for all $v\in\ell^{2}(\mathbb{N})$. Letting $U_{\mathcal{H}}$ denote the unitary
$\frac{\sqrt{2}}{2}\begin{pmatrix}I_{\mathcal{H}}&-iI_{\mathcal{H}}\\ -I_{\mathcal{H}}&I_{\mathcal{H}}\end{pmatrix}\in B(\ell^{2}({\mathbb{N}})\oplus\ell^{2}({\mathbb{N}})),$
we find that each element of the compact operator tuple $U_{\mathcal{H}}^{*}(A\oplus\overline{A})U_{\mathcal{H}}$ maps the real part of $\ell^{2}({\mathbb{N}})\oplus\ell^{2}({\mathbb{N}})$ to the real part of $\ell^{2}({\mathbb{N}})\oplus\ell^{2}({\mathbb{N}})$. Therefore, we can treat $U_{\mathcal{H}}^{*}(A\oplus\overline{A})U_{\mathcal{H}}$ as a tuple of bounded compact operators on $\ell^{2}({\mathbb{N}})\oplus\ell^{2}({\mathbb{N}})$ viewed a Hilbert space over $\mathbb{R}$. From this we obtain that ${\mathcal{D}}_{A}^{\mathbb{R}}={\mathcal{D}}_{U_{\mathcal{H}}^{*}(A\oplus\overline{A})U_{\mathcal{H}}}^{\mathbb{R}}$ is a real generalized free spectrahedron. ∎
8.3. Proof of Lemma 5.4
Proof.
We first prove Item (1). When working with a free spectrahedrop or a generalized free spectrahedron, one can apply Lemmas 3.1 and 4.4 together with the strategy used to prove [24, Lemma 3.2 (1)]. However, an alternative approach is needed for general matrix convex sets. For this case, Let $X\in K(n)$ and $Y\in K(m)$ and let $\beta\in{\mathfrak{D}_{X}(K)}$ and $\eta\in{\mathfrak{D}_{Y}(K)}$. Then using Lemma 2.2 there exist constants $c_{1},c_{2}>0$ and tuples $\gamma,\sigma\in\mathbb{R}^{g}$ such that
$\begin{pmatrix}X&0&c_{1}\beta\\ 0&Y&0\\ c_{1}\beta^{*}&0&\gamma\end{pmatrix}\in K\qquad\mathrm{and}\qquad\begin{pmatrix}X&0&0\\ 0&Y&c_{2}\eta\\ 0&c_{2}\eta^{*}&\sigma\end{pmatrix}\in K.$
It follows that
$\begin{pmatrix}\beta^{*}&0\end{pmatrix}^{*},\begin{pmatrix}0&\eta^{*}\end{pmatrix}^{*}\in{\mathfrak{D}_{K}(X\oplus Y)},$
from which we obtain
${\mathfrak{D}_{X\oplus Y}(K)}\supseteq\left\{\left(\beta^{*}\ \eta^{*}\right)^{*}\in M_{(n+m)\times 1}(\mathbb{F})^{g}|\ \beta\in{\mathfrak{D}_{X}(K)}\mathrm{\ and\ }\eta\in{\mathfrak{D}_{Y}(K)}\right\}.$
For the reverse containment, suppose $\left(\beta^{*}\ \eta^{*}\right)^{*}\in{\mathfrak{D}_{K}(X\oplus Y)}$. Again using Lemma 2.2 shows there is a constant $c>0$ and a tuple $\gamma\in\mathbb{R}^{g}$ such that
$\begin{pmatrix}X&0&c\beta\\ 0&Y&c\eta\\ c\beta^{*}&c\eta^{*}&\sigma\end{pmatrix}\in K.$
Using the matrix convexity of $K$, we obtain
$\begin{pmatrix}1&0&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}X&0&c\beta\\ 0&Y&c\eta\\ c\beta^{*}&c\eta^{*}&\gamma\end{pmatrix}\begin{pmatrix}1&0\\ 0&0\\ 0&1\end{pmatrix}=\begin{pmatrix}X&c\beta\\ c\beta^{*}&\gamma\end{pmatrix}\in K,$
and similarly
$\begin{pmatrix}0&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}X&0&c\beta\\ 0&Y&c\eta\\ c\beta^{*}&c\eta^{*}&\gamma\end{pmatrix}\begin{pmatrix}0&0\\ 1&0\\ 0&1\end{pmatrix}=\begin{pmatrix}X&c\eta\\ c\eta^{*}&\gamma\end{pmatrix}\in K.$
We conclude $\beta\in{\mathfrak{D}_{X}(K)}$ and $\eta\in{\mathfrak{D}_{Y}(K)}$, which proves the reverse containment.
The proof of Items (2) and (3) use the same approach of [24, Lemma 3.2 (2) and (3)] together with an application of Lemma 2.2. In particular, for a unitary $U\in M_{n}(\mathbb{F})$, we have
$\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K\iff\begin{pmatrix}U^{*}XU&U^{*}\beta\\ \beta^{*}U&\gamma\end{pmatrix}=\begin{pmatrix}U^{*}&0\\ 0&1\end{pmatrix}\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\begin{pmatrix}U&0\\ 0&1\end{pmatrix}\in K$
Similarly, if $K$ is closed under complex conjugation, then we immediately have
$\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in K\iff\begin{pmatrix}\overline{X}&\overline{\beta}\\ \overline{\beta^{*}}&\overline{\gamma}\end{pmatrix}\in K.$
Using the above with Lemma 2.2 completes the proof of Items (2) and (3).
Item (4) is an immediate consequence of Items (1), (2), and (3). ∎